### v1

# Robust Inference for Non-Linear Regression Models with Applications in Enzyme Kinetics

Abstract
Despite linear regression being the most popular statistical modelling technique, in real-life we often need to deal with situations where the true relationship between the response and the covariates is nonlinear in parameters. In such cases one needs to adopt appropriate non-linear regression (NLR) analysis, having wider applications in biochemical and medical studies among many others. In this paper we propose a new improved robust estimation and testing methodologies for general NLR models based on the minimum density power divergence approach and apply our proposal to analyze the widely popular Michaelis-Menten (MM) model in enzyme kinetics. We establish the asymptotic properties of our proposed estimator and tests, along with their theoretical robustness characteristics through influence function analysis. For the particular MM model, we have further empirically justified the robustness and the efficiency of our proposed estimator and the testing procedure through extensive simulation studies and several interesting real data examples of enzyme-catalyzed (biochemical) reactions.
Keywords: Robustness, density power divergence, robust Wald-type tests, Michaelis-Menten equation
1 Introduction
Linear regression is arguably the most used statistical modeling technique to investigate relationships between variables. However, even for normally distributed continuous response variables, the linearity assumption appears to be restrictive in several real-life situations where the true relationship between the response and the covariates may be best modelled by a suitable non-linear function in parameters. Such scenarios frequently arise in biomedical and biochemical researches, besides statistical signal analyses and many other domains. For example, a popular such non-linear relationship in enzyme kinetics is known as the Michaelis-Menten (MM) equation. This MM equation, originally developed by Leonor Michaelis and Maud Menten in 1913 [30], has now become the foundational basis of enzymology for studying the rates of enzyme catalyzed chemical reactions; see, e.g., [2, 3, 22, 31, 35]. For such a reaction, if the reaction velocity is $v$ at a substrate concentration $s$, then the MM equation says
$v=\frac{V_{max}s}{K_{m}+s},$ (1)
where $K_{m}$ is the pseudo-equilibrium constant of the reaction, called Michaelis constant, and $V_{max}$ is the maximum velocity attainable (theoretically) at an infinite substrate concentration saturated with the enzyme. Here $K_{m}$ and $V_{max}$ are the two reaction-specific unknown (model) parameters which are estimated from experimental data, consisting of the observed reaction velocities ($v_{i}$) at different (often pre-fixed) substrate concentrations ($s_{i}$), for a specific chemical reaction, assuming additive (and independent) Gaussian error in each observation $v_{i}$ (for $i=1,\ldots,n$, say). This estimation setup can be expressed as a non-linear regression (NLR) model given by
$v_{i}=\mu(s_{i},\bm{\beta})+\epsilon_{i},~{}~{}~{}~{}i=1,\ldots,n,$ (2)
where $\mu(s_{i},\bm{\beta})=\frac{\beta_{1}s_{i}}{\beta_{2}+s_{i}}$ and $\bm{\beta}=(\beta_{1},\beta_{2})^{T}=(V_{max},K_{m})^{T}$. The estimation of the model parameter $\bm{\beta}$ and the subsequent inference are commonly performed using the non-linear least-squares or the maximum likelihood approach [18, 37, 21]. These inferences are also extended to variants of the MM model for specific applications; see, e.g., [38, 12]. The MM model (1) can also be analyzed using a linearizing transformation, called Woolf transformation, used by Currie [11]. However, for that case it is not possible to take additive errors, as in model (2).
More generally, given $n$ observations $y_{i}$ and $\bm{x}_{i}$ on a (real valued) response variable $Y$ and associated $k$-dimensional covariates $\bm{X}$, respectively, for $i=1,\ldots,n$, a typical NLR model can be expressed as [9, 34]
$y_{i}=\mu(\bm{x_{i}},\bm{\beta})+\epsilon_{i},\ i=1,2,\ldots,n,$ (3)
where $\bm{\beta}$ is a $p$-dimensional regression parameter vector and $\epsilon_{i}$’s are independently and identically distributed (IID) random errors with mean 00 and variance $\sigma^{2}$. Here the form of the mean (or regression) function $\mu(\cdot,\cdot)$ is assumed to be known (pre-specified), continuous, and non-linear in the parameter $\bm{\beta}$. The full parameter vector in this NLR model is then given by $\bm{\theta}=(\bm{\beta}^{T},\sigma^{2})^{T}\in\bm{\Theta}\subseteq\mathbb{R}^{p}\times\mathbb{R}^{+}$. It is to be noted that, since $\mu(\bm{x_{i}},\bm{\beta})$ is non-linear in $\bm{\beta}$, dimensions of $\bm{x_{i}}$ and $\bm{\beta}$ are not necessarily the same (i.e., $p$ may or may not be equal to $k$). There is a vast amount of literature on the classical ordinary least-square (OLS) and the maximum likelihood (ML) methods under NLR models. Asymptotic properties of the least-square estimator under the NLR model were initially studied by Jennrich [20] and Wu [40]. It can also be easily observed that, under the normality assumption on the IID errors $\epsilon_{i}$’s in (3), both the OLS estimator and the ML estimator (MLE) of the parameter $\bm{\beta}$ coincide. Discussions on the MLE under the NLR model (3) and its asymptotic properties are available in, e.g., Seber and Wild [34]. Further refinements and advancements in both theory and computation of these estimators under general NLR models and several of its important special cases are available in, e.g., [32, 23, 24, 25], among many others.
However, the main drawback of the least-squares or the maximum likelihood based inference is their high sensitivity to the presence of outliers and other contaminations in the observed data, even in a small proportion. Since we frequently encounter outliers and influential observations in real-life datasets, there is a clear need for suitable robust estimation and inference procedures yielding stable and accurate results even under data contamination. But, although there exists a vast literature on robust statistical procedures for linear and generalized linear regression models, there have been comparatively fewer works developing robust procedures for NLR models. In particular, Liu et al. [27] and Marasovic et al. [28] studied robust M-estimators under the NLR models with Huber’s and Tukey’s loss/weight functions, respectively. Tabatabai et al. [36] proposed a modified M-estimation procedure (to be referred to as the KPS estimator) whereas Liu et al. [26] studied the median of means (MOM) estimator under the NLR model. Giradi et al. [17] used Tsallis score for robust estimation under a particular example of 3-parameter log-logistic NLR model, for analyzing COVID-19 contagion data of Italy, but they avoided many mathematical details and justifications of the estimation procedure. Further, all these existing works only dealt with the estimation problem and there is no concrete work on robust hypotheses testing available in the literature of NLR models. In the context of MM models in enzyme kinetics, only the KPS and the M-estimators are explored to generate robust solutions under possible data contaminations; see, e.g., [36, 28, 27], among others.
In this paper our aim is to develop an improved robust estimator and associated robust testing procedure for the general class of NLR models based on the popular minimum density power divergence (DPD) approach of Basu et al. [6], and apply them for robust analysis of MM models with greater accuracy compared to the existing alternatives. Due to various nice properties, the DPD based approach has been successfully implemented in several statistical problems, including regressions; see, e.g., [14, 15, 16, 7, 5, 13, 10, 4]. For the sake of completeness, a brief background about the DPD and the minimum DPD estimator (MDPDE) is provided in Appendix A.1. Here we will extend the definition of the MDPDE for the general NLR models and discuss its asymptotic properties (consistency and asymptotic normality) under simple verifiable conditions. We will justify the robustness property of our proposed estimator by examining the boundedness of its influence function. We will additionally develop a robust hypothesis testing procedure for general NLR models based on our proposed estimator (i.e., the MDPDE) and examine the influence functions of our test statistics to justify their claimed robustness. We will then apply our proposed methodologies for robust inference under the MM model (2), with simplified theoretical properties and extensive empirical illustrations of their finite-sample performances. In particular, we will develop and study a robust test procedure for testing the validity of the MM model for a given (possibly contaminated) experimental dataset. Finally, we will present some interesting real data examples of different enzyme kinetic reactions.
The rest of the paper is organized as follows: In Section 2, we present our proposed MDPDE along with the derivation of its asymptotic and robustness properties. In Section 3 we have discussed the proposed robust hypothesis testing procedures utilizing the MDPDEs and studied their asymptotic properties and their influence functions. Section 4 is devoted to the application of our proposed methods to the Michaelis-Menten model (2) along with extensive simulation studies illustrating their improved finite-sample performances in comparison to the existing robust approaches. Section 5 consists of discussion on some data driven procedures for finding an optimal choice of the MDPDE tuning parameter. Section 6 presents analyses of some real data examples and final concluding remarks are given in Section 7. For brevity in presentation, some background details, proofs of the results, and some numerical illustrations are provided in Appendices A-C.
2 Robust parameter estimation under general NLR models
2.1 The MDPDE
Let us consider the NLR model given in (3) with the error assumption $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, independently, for all $i=1,2,\ldots,n$. So, we have $n$ independent and non-homogeneous (INH) observations $y_{1},y_{2},\ldots,y_{n}$ with
$y_{i}\sim\mathcal{N}\left(\mu(\bm{x_{i}},\bm{\beta}),\sigma^{2}\right),\mbox{ independently, for }i=1,\ldots,n.$
So, we can follow the general theory of the MDPDE under INH setup, as described briefly in Appendix A.1 following Ghosh and Basu [14], to define and study the MDPDE of the parameters under the present NLR model. Continuing with the notations from Appendix A.1, in this case, we have the model density for $i$-th observation as $f_{i}(\cdot;\bm{\theta})\equiv\mathcal{N}\left(\mu(\bm{x_{i}},\bm{\beta}),\sigma^{2}\right)$, so that we get
$V_{i}(y_{i};\bm{\theta})=\frac{1}{\left(\sigma\sqrt{2\pi}\right)^{\alpha}\sqrt{1+\alpha}}-\frac{1+\alpha}{\alpha}\frac{1}{\left(\sigma\sqrt{2\pi}\right)^{\alpha}}e^{-\frac{\alpha}{2\sigma^{2}}(y_{i}-\mu(\bm{x_{i}},\bm{\beta}))^{2}},\ \ \alpha>0.$ (4)
Then the MDPDE of the parameter $\bm{\theta}=\left(\bm{\beta}^{T},\sigma^{2}\right)^{T}$ under our NLR model can be defined as the minimizer of the objective function
$H_{n}(\bm{\theta})=\frac{1}{\left(\sigma\sqrt{2\pi}\right)^{\alpha}\sqrt{1+\alpha}}-\frac{1+\alpha}{\alpha}\frac{1}{\left(\sigma\sqrt{2\pi}\right)^{\alpha}}\frac{1}{n}\sum_{i=1}^{n}e^{-\frac{\alpha}{2\sigma^{2}}(y_{i}-\mu(\bm{x_{i}},\bm{\beta}))^{2}},\ \ \alpha>0.$ (5)
Now, differentiating (5) partially with respect to $\bm{\beta}$ and $\sigma^{2}$, and equating to 00, we get the estimating equations for the MDPDE as given by
$\sum_{i=1}^{n}e^{-\frac{\alpha}{2\sigma^{2}}(y_{i}-\mu(\bm{x_{i}},\bm{\beta}))^{2}}(y_{i}-\mu(\bm{x_{i}},\bm{\beta}))\nabla_{\bm{\beta}}\mu(\bm{x_{i}},\bm{\beta})=0,$ (6)
$\sum_{i=1}^{n}e^{-\frac{\alpha}{2\sigma^{2}}(y_{i}-\mu(\bm{x_{i}},\bm{\beta}))^{2}}\left[1-\frac{(y_{i}-\mu(\bm{x_{i}},\bm{\beta}))^{2}}{\sigma^{2}}\right]=\frac{n\alpha}{(1+\alpha)^{3/2}},$ (7)
where $\nabla_{\bm{\beta}}$ denotes the first order partial derivative with respect to $\bm{\beta}$. To obtain the MDPDE $\bm{\widehat{\theta}_{\alpha}}=\left(\bm{\widehat{\beta}_{\alpha}}^{T},\widehat{\sigma}_{\alpha}^{2}\right)^{T}$ with tuning parameter $\alpha>0$, in practice, one can numerically solve the above two estimating equations simultaneously for $\bm{\theta}=\left(\bm{\beta}^{T},\sigma^{2}\right)^{T}$, which is equivalent to minimize (5). Note that, the MDPDE coincides with the MLE at $\alpha=0$ in a limiting sense, and provides a robust generalization of the MLE at $\alpha>0$.
Note that, an important advantage of our MDPDE over its existing robust competitors (briefly described in Appendix A.2) is that the MDPDEs additionally yield a robust estimate of the scale parameter $\sigma$ simultaneously with the estimation of the regression parameter $\bm{\beta}$. Further, for the MDPDE, the problem with multiple roots of the estimation equations can be easily solved using the associated objective function in (5).
2.2 Asymptotic properties
For simplicity, let us assume that the true data generating density $g_{i}$ of $y_{i}$ belongs to the corresponding parametric model family
$\mathcal{F}_{i}=\left\{\mathcal{N}\left(\mu(\bm{x_{i}},\bm{\beta}),\sigma^{2}\right):\bm{\theta}=\left(\bm{\beta}^{T},\sigma^{2}\right)^{T}\in\bm{\Theta}\right\},~{}~{}~{}i=1,2,\ldots,n,$ (8)
i.e., $g_{i}\equiv\mathcal{N}\left(\mu(\bm{x_{i},\beta_{0}}),\sigma_{0}^{2}\right),\mbox{ for some }\bm{\theta_{0}}=\left(\bm{\beta_{0}}^{T},\sigma_{0}^{2}\right)^{T}\in\bm{\Theta}$. Our NLR model (3) can be expressed in vectorized form as
$\bm{y}=\bm{\mu(\beta)}+\bm{\epsilon},$ (9)
where $\bm{y}=(y_{1},\ldots,y_{n})^{T},\bm{\epsilon}=(\epsilon_{1},\ldots,\epsilon_{n})^{T}$ and $\bm{\mu(\beta)}=\left(\mu(\bm{x_{1},\beta}),\ldots,\mu(\bm{x_{n},\beta})\right)^{T}$. Now we define the $n\times p$ matrix
$\bm{\dot{\mu}(\beta)}=\frac{\partial}{\partial\bm{\beta}}\bm{\mu(\beta)}=\begin{pmatrix}\nabla_{\bm{\beta}}\mu(\bm{x_{1}},\bm{\beta})^{T}\\ \nabla_{\bm{\beta}}\mu(\bm{x_{2}},\bm{\beta})^{T}\\ \vdots\\ \nabla_{\bm{\beta}}\mu(\bm{x_{n}},\bm{\beta})^{T}\\ \end{pmatrix},$ (10)
and the $(p+1)\times(p+1)$ matrices
$\bm{\Psi}_{n}=\begin{bmatrix}\frac{\zeta_{\alpha}}{n}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)&\bm{0}_{p}\\ \\ \bm{0}_{p}^{T}&\varsigma_{\alpha}\end{bmatrix}\mbox{ and }\bm{\Omega}_{n}=\begin{bmatrix}\frac{\zeta_{2\alpha}}{n}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)&\bm{0}_{p}\\ \\ \bm{0}_{p}^{T}&\varsigma_{2\alpha}-\frac{\alpha^{2}}{4}\zeta_{\alpha}^{2}\end{bmatrix},$ (11)
where
$\zeta_{\alpha}=(2\pi)^{-\alpha/2}\sigma^{-(\alpha+2)}(1+\alpha)^{-3/2}\mbox{ and }\varsigma_{\alpha}=(2\pi)^{-\alpha/2}\sigma^{-(\alpha+4)}\frac{2+\alpha^{2}}{4(1+\alpha)^{5/2}}.$
We now present some conditions on the NLR model, under which the asymptotic properties of the MDPDE $\bm{\widehat{\theta}_{\alpha}}$ would be derived. Let us denote $\bm{\Theta_{\beta}}\ (\subseteq\mathbb{R}^{p})$ be the parameter space of the model parameter $\bm{\beta}$ of the NLR model (3) and $\sigma^{2}>0$, so that $\bm{\Theta}=\bm{\Theta_{\beta}}\times\mathbb{R}^{+}$.
Conditions for consistency and asymptotic normality:
(R1) Let $(\bm{\dot{\mu}})_{ij}$ be the $(i,j)$-th element of the matrix $\bm{\dot{\mu}(\beta)}$, given in (10). Then, for all $j,k,l$ and for all $\bm{\beta}\in\omega$, an open subset of $\bm{\Theta_{\beta}}$, the mean function $\mu(\bm{x_{i}},\bm{\beta})$ is thrice continuously differentiable with respect to $\bm{\beta}$ with
$\sup_{n>1}\max_{1\leq i\leq n}\left|(\bm{\dot{\mu}})_{ij}\right|=O(1),\ \ \ \ \sup_{n>1}\max_{1\leq i\leq n}\left|(\bm{\dot{\mu}})_{ij}(\bm{\dot{\mu}})_{ik}\right|=O(1),$ (12)
and
$\frac{1}{n}\sum_{i=1}^{n}\left|(\bm{\dot{\mu}})_{ij}(\bm{\dot{\mu}})_{ik}(\bm{\dot{\mu}})_{il}\right|=O(1).$ (13)
(R2) The matrix $\bm{\dot{\mu}(\beta)}$ satisfies
$\inf_{n}\left[\mbox{minimum eigenvalue of }\frac{1}{n}\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right]>0,$ (14)
which says that $\bm{\dot{\mu}(\beta)}$ is of full column rank, and
$n\max_{1\leq i\leq n}\left[\nabla_{\bm{\beta}}\mu(\bm{x_{i}},\bm{\beta})^{T}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)^{-1}\nabla_{\bm{\beta}}\mu(\bm{x_{i}},\bm{\beta})\right]=O(1).$ (15)
With the above mentioned conditions, the asymptotic distributions of the MDPDEs of the parameters of the NLR model (3) are given in the following theorem; its proof is given in Appendix B.1.
Theorem 2.1.
Under the setup of the NLR model in (3), suppose that the true data generating distributions belong to the corresponding parametric model families with the true parameter value being $\bm{\theta}_{0}=\left(\bm{\beta}_{0}^{T},\sigma_{0}^{2}\right)^{T}$ and Conditions (R1)-(R2) hold. Then, for any fixed $\alpha\geq 0$, we have the following results.
  1. (i)
    There exists a consistent sequence of roots $\bm{\widehat{\theta}_{\alpha}}=\left(\bm{\widehat{\beta}_{\alpha}}^{T},\widehat{\sigma}_{\alpha}^{2}\right)^{T}$ of the MDPDE estimating equations (6) and (7).
  2. (ii)
    Asymptotic distributions of $\bm{\widehat{\beta}_{\alpha}}$ and $\widehat{\sigma}_{\alpha}^{2}$ are independent.
  3. (iii)
    Asymptotically, as $n\to\infty$,
    $\left(\bm{\dot{\mu}(\beta_{0})}^{T}\bm{\dot{\mu}(\beta_{0})}\right)^{1/2}\left(\bm{\widehat{\beta}_{\alpha}}-\bm{\beta_{0}}\right)\xrightarrow{\mathcal{D}}\mathcal{N}_{p}\left(\bm{0}_{p},\ v_{1\alpha}\sigma_{0}^{2}\mathbb{I}_{p}\right),$
    and
    $\sqrt{n}\left(\widehat{\sigma}_{\alpha}^{2}-\sigma_{0}^{2}\right)\xrightarrow{\mathcal{D}}\mathcal{N}\left(0,\ v_{2\alpha}\sigma_{0}^{4}\right),$
where
$v_{1\alpha}=\frac{1}{\sigma^{2}}\frac{\zeta_{2\alpha}}{\zeta_{\alpha}^{2}}=\left(1+\frac{\alpha^{2}}{1+2\alpha}\right)^{3/2},$
$v_{2\alpha}=\frac{1}{\sigma^{4}}\frac{\varsigma_{2\alpha}-\frac{\alpha^{2}}{4}\zeta_{\alpha}^{2}}{\varsigma_{\alpha}^{2}}=\frac{4}{\left(2+\alpha^{2}\right)^{2}}\left[2\left(1+2\alpha^{2}\right)\left(1+\frac{\alpha^{2}}{1+2\alpha}\right)^{5/2}-\alpha^{2}(1+\alpha)^{2}\right].$
Remark 2.1.
For $\alpha=0$, the above theorem gives the asymptotic normality result for the OLS or the MLE under NLR models (with IID normal errors), which is exactly the same as discussed in Theorem 5 of Wu [40], for the OLS.
Let us now study the asymptotic relative efficiency (ARE) of the MDPDE with respect to the fully efficient MLE. From Theorem 2.1 it is easy to see that the ARE of the MDPDE $\bm{\widehat{\beta}_{\alpha}}$ of $\bm{\beta}$ is the same for all $\beta_{i}$’s and is given by
$\frac{v_{10}}{v_{1\alpha}}\times 100\%=\left(1+\frac{\alpha^{2}}{1+2\alpha}\right)^{-3/2}\times 100\%.$
Similarly the ARE of $\widehat{\sigma}_{\alpha}^{2}$ is given by
$\frac{v_{20}}{v_{2\alpha}}\times 100\%=\frac{(2+\alpha^{2})^{2}}{2}\left[2(1+2\alpha^{2})\left(1+\frac{\alpha^{2}}{1+2\alpha}\right)^{5/2}-\alpha^{2}(1+\alpha)^{2}\right]^{-1}\times 100\%.$
Note that, these AREs are independent of the choice of the (non-linear) regression function $\mu(\cdot,\cdot)$ and exactly the same as those obtained for the linear regression model in Ghosh and Basu [14]. The exact values of the ARE of the MDPDE of $\bm{\beta}$ for different $\alpha$ are reported in Table 1, along with the same for the existing robust competitors (as described in Appendix A.2). From the table, it is evident that for the MDPDE, the loss in efficiency is very small for smaller values of $\alpha$, in comparisons to other competing estimators at suitable tuning parameter values expecting to produce similar robustness properties.
Table 1: AREs of different estimators of $\bm{\beta}$ of the NLR model at several values of the associated tuning parameters ($c$, $w$ or $\alpha$)
Huber’s M-est Tukey’s M-est KPS est MDPDE
c ARE c ARE w ARE $\alpha$ ARE
1.000 0.466 3.000 0.776 0.5 0.911 0.1 0.988
1.345 0.589 3.500 0.858 1.0 0.647 0.3 0.921
1.700 0.710 4.000 0.910 1.3 0.503 0.5 0.838
2.000 0.802 4.685 0.950 1.5 0.424 0.7 0.757
2.500 0.916 5.500 0.973 1.7 0.357 1.0 0.650
2.3 Influence function and sensitivity analysis
Influence function (IF) is a measure of local robustness property of an estimator. Here, we will follow the general theory from Ghosh and Basu [14], as presented briefly in Appendix A.1.
Firstly, the minimum DPD functional $\bm{T_{\alpha}}=\left(\bm{T_{\alpha}^{\beta}},T_{\alpha}^{\sigma^{2}}\right)$ of the parameter $\bm{\theta}=\left(\bm{\beta}^{T},\sigma^{2}\right)^{T}$ under the setup of the NLR model (3), with true distribution $g_{i}$ and $f_{i}(\cdot;\bm{\theta})\in\mathcal{F}_{i}$, is defined by
[…]
where $\bm{G}=(G_{1},\cdots,G_{n})^{T}$ with $G_{i}$ being the distribution function for the density $g_{i}$ for each $i=1,\ldots,n$. Then, assuming the true distributions belonging to the corresponding model families and using the general expression of IF under INH setup (as given in Eq. (38) in Appendix A.1), the IF of $\bm{T}_{\alpha}$ with contamination only in the $i_{0}$-th observation at the contamination point $t_{i_{0}}$ is obtained as
$IF_{i_{0}}(t_{i_{0}},\bm{T}_{\alpha},\bm{F}_{\bm{\theta}})=\begin{pmatrix}\frac{1}{\zeta_{\alpha}}f_{i_{0}}(t_{i_{0}};\bm{\theta})^{\alpha}\frac{(t_{i_{0}}-\mu(\bm{x_{i_{0}}},\bm{\beta}))}{\sigma^{2}}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)^{-1}\nabla_{\bm{\beta}}\mu(\bm{x_{i_{0}}},\bm{\beta})\\ \\ \frac{1}{n\varsigma_{\alpha}}\left[\left\{\frac{(t_{i_{0}}-\mu(\bm{x_{i_{0}}},\bm{\beta}))^{2}}{2\sigma^{4}}-\frac{1}{2\sigma^{2}}\right\}f_{i_{0}}(t_{i_{0}};\bm{\theta})^{\alpha}+\frac{\alpha}{2}\zeta_{\alpha}\right]\end{pmatrix},$
where $\bm{F}_{\bm{\theta}}=\left(F_{1,\bm{\theta}},F_{2,\bm{\theta}},\ldots,F_{n,\bm{\theta}}\right)^{T}$ with $F_{i,\bm{\theta}}$ being the distribution function for the model density $f_{i}(\cdot;\bm{\theta})$. On simplification, the IFs of $\bm{T}_{\alpha}^{\bm{\beta}}$ and $T_{\alpha}^{\sigma^{2}}$ with contamination only in the $i_{0}$-th observations at the contamination point $t_{i_{0}}$, respectively, become
$IF_{i_{0}}(t_{i_{0}},\bm{T}_{\alpha}^{\bm{\beta}},\bm{F}_{\bm{\theta}})=\\ (1+\alpha)^{\frac{3}{2}}(t_{i_{0}}-\mu(\bm{x_{i_{0}}},\bm{\beta}))e^{-\frac{\alpha}{2\sigma^{2}}(t_{i_{0}}-\mu(\bm{x_{i_{0}}},\bm{\beta}))^{2}}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)^{-1}\nabla_{\bm{\beta}}\mu(\bm{x_{i_{0}}},\bm{\beta}),$ (17)
and
$IF_{i_{0}}(t_{i_{0}},T_{\alpha}^{\sigma^{2}},\bm{F}_{\bm{\theta}})=\\ \frac{2(1+\alpha)^{\frac{5}{2}}}{n(2+\alpha^{2})}\left\{(t_{i_{0}}-\mu(\bm{x_{i_{0}}},\bm{\beta}))^{2}-\sigma^{2}\right\}e^{-\frac{\alpha}{2\sigma^{2}}(t_{i_{0}}-\mu(\bm{x_{i_{0}}},\bm{\beta}))^{2}}+\frac{2\alpha(1+\alpha)^{2}}{n(2+\alpha^{2})}.$ (18)
Note that, the functions $z\mapsto e^{-kz^{2}},\ z\mapsto ze^{-kz^{2}},\ z\mapsto z^{2}e^{-kz^{2}}$ are all bounded in $\mathbb{R}$, for any $k>0$, and hence the above IFs in (17), (18) are bounded in the contamination point $t_{i_{0}},\mbox{ for all }\alpha>0\mbox{ and any }i_{0}$. This ensures that, for $\alpha>0$, the MDPDEs are robust with respect to any outlying observation. Also note that, the IFs are unbounded at $\alpha=0$, which is actually the case of the non-robust MLEs.
Further, according to the general formula given in Eq. (39) in Appendix A.1, the IFs of $\bm{T^{\beta}}_{\alpha}$ and $T_{\alpha}^{\sigma^{2}}$ with contamination in all the observations at the contamination points in $\bm{t}=(t_{1},\cdots,t_{n})^{T}$, are, respectively, given by
$IF(\bm{t},\bm{T}_{\alpha}^{\bm{\beta}},\bm{F}_{\bm{\theta}})=\\ (1+\alpha)^{\frac{3}{2}}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)^{-1}\sum_{i=1}^{n}\nabla_{\bm{\beta}}\mu(\bm{x_{i}},\bm{\beta})(t_{i}-\mu(\bm{x_{i}},\bm{\beta}))e^{-\frac{\alpha}{2\sigma^{2}}(t_{i}-\mu(\bm{x_{i}},\bm{\beta}))^{2}},$ (19)
$IF(\bm{t},T_{\alpha}^{\sigma^{2}},\bm{F}_{\bm{\theta}})=\\ \frac{2(1+\alpha)^{\frac{5}{2}}}{n(2+\alpha^{2})}\sum_{i=1}^{n}\left\{(t_{i}-\mu(\bm{x_{i}},\bm{\beta}))^{2}-\sigma^{2}\right\}e^{-\frac{\alpha}{2\sigma^{2}}(t_{i}-\mu(\bm{x_{i}},\bm{\beta}))^{2}}+\frac{2\alpha(1+\alpha)^{2}}{(2+\alpha^{2})}.$ (20)
It is easy to observe that these IFs are also bounded in each argument $t_{i},\mbox{ for }i=1,2,\ldots,n,\mbox{ for all }\alpha>0$, and unbounded at $\alpha=0$, as before.
Let us now consider some IF based summary measures of robustness for our NLR model. The gross error sensitivity of the minimum DPD functional $\bm{T^{\beta}}_{\alpha}$ of $\bm{\beta}$ at the vector of model distributions $\bm{F_{\theta}}$ with contamination only in the $i_{0}$-th observation is given by
$\gamma_{i_{0}}^{u}(\bm{T}_{\alpha}^{\bm{\beta}},\bm{F}_{\bm{\theta}})=\sup_{t}\left\{\lVert IF_{i_{0}}(t,\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})\rVert\right\}\\ =\left\{\begin{matrix}\frac{(1+\alpha)^{\frac{3}{2}}}{\sqrt{\alpha}}\sigma e^{-\frac{1}{2}}\lVert\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)^{-1}\nabla_{\bm{\beta}}\mu(\bm{x_{i_{0}}},\bm{\beta})\rVert,&\mbox{if}&\alpha>0,\\ \infty,&\mbox{if}&\alpha=0.\end{matrix}\right.$ (21)
But this measure does not satisfy the scale invariance property with respect to scale transformation of the components of $\bm{\beta}$. To overcome this problem, we consider a modified measure, namely the self standardized sensitivity. For contamination only in the $i_{0}$-th observation, it is given by
$\gamma_{i_{0}}^{s}(\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})=\sup_{t}\left\{IF_{i_{0}}(t,\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})^{T}(\bm{\Psi}_{pn}^{-1}\bm{\Omega}_{pn}\bm{\Psi}_{pn}^{-1})^{-1}IF_{i_{0}}(t,\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})\right\}^{1/2}\\ =\left\{\begin{matrix}\frac{(1+2\alpha)^{\frac{3}{4}}}{\sqrt{n\alpha}}e^{-\frac{1}{2}}\left\{\nabla_{\bm{\beta}}\mu(\bm{x_{i_{0}},\beta})^{T}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)^{-1}\nabla_{\bm{\beta}}\mu(\bm{x_{i_{0}},\beta})\right\}^{1/2},&\mbox{if}&\alpha>0,\\ \infty,&\mbox{if}&\alpha=0,\end{matrix}\right.$ (22)
where $\bm{\Psi}_{pn},\bm{\Omega}_{pn}$ are the $p\times p$ principal diagonal blocks of the matrices $\bm{\Psi}_{n},\bm{\Omega}_{n}$, respectively. Note that, both the above sensitivity measures are decreasing function in $\alpha\ (>0)$. So, it is evident that, with the presence of outliers, the robustness of the MDPDE of $\bm{\beta}$ increases with increasing values of $\alpha$. However, in Section 2.2 we have observed that the ARE of MDPDE of $\bm{\beta}$ (under pure data) decreases with increasing values of $\alpha$. Together, the tuning parameter $\alpha$ acts as a trade-off between efficiency and robustness of the MDPDE under NLR models, in consistence with the existing literature of DPD based procedures.
For contamination in all or some observations, the gross error sensitivity and the self standardized sensitivity can similarly be defined as in (21) and (22), but replacing $IF_{i_{0}}(t,\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})$ by $IF(\bm{t},\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})$ and taking supremum over all possible $t_{1}\ldots,t_{n}$. The resulting implications are again the same, and hence these details are omitted for brevity.
3 Robust Wald-type tests based on the MDPDE
Basu et al. [7] first defined and studied a generalized class of robust Wald-type tests based on the MDPDE under the IID setup, which was later extended for the general INH setup in Basu et al. [5]. Here we will use the same approach to develop robust tests of hypotheses for our NLR models.
3.1 General Wald-type tests of NLR model parameters
Consider the setup of the NLR model (3) and suppose that we wish to test the general linear hypothesis on the regression parameters $\bm{\beta}$, as given by
$H_{0}:\bm{L}\bm{\beta}=\bm{l}_{0}\mbox{ against }H_{1}:\bm{L\beta}\neq\bm{l}_{0},$ (23)
where $\bm{L}$ is a given $r\times p$ matrix with rank$(\bm{L})=r$ and $\bm{l}_{0}$ is a given $r$-vector. Recall that the full parameter vector is $\bm{\theta}=\left(\bm{\beta}^{T},\sigma^{2}\right)^{T}$ with the parameter space $\bm{\Theta}\subseteq\mathbb{R}^{p}\times\mathbb{R}^{+}$. The restricted parameter space under the null hypothesis in (23) is given by
$\bm{\Theta_{0}}=\left\{\left(\bm{\beta}^{T},\sigma^{2}\right)^{T}:\bm{L\beta=l_{0}}\mbox{ and }0<\sigma^{2}<\infty\right\}.$
Also, the MDPDE of $\bm{\theta}=\left(\bm{\beta}^{T},\sigma^{2}\right)^{T}$ with tuning parameter $\alpha\geq 0$ has been denoted by $\bm{\widehat{\theta}_{\alpha}}=\left(\bm{\widehat{\beta}_{\alpha}}^{T},\widehat{\sigma}_{\alpha}^{2}\right)^{T}$. Then, we define the corresponding Wald-type test statistics for testing (23) as given by
$W_{n,\alpha}=\frac{1}{v_{1\alpha}\widehat{\sigma}_{\alpha}^{2}}\bm{\left(L\widehat{\beta}_{\alpha}-l_{0}\right)^{T}\left[L\left(\dot{\mu}\left(\widehat{\beta}_{\alpha}\right)^{T}\dot{\mu}\left(\widehat{\beta}_{\alpha}\right)\right)^{-1}L^{T}\right]^{-1}\left(L\widehat{\beta}_{\alpha}-l_{0}\right)}.$
Note that, under the null hypothesis in (23), one can easily show that, $W_{n,\alpha}$ asymptotically follows $\chi^{2}_{r}$, the chi-square distribution with $r$ degrees of freedom. So, the null hypothesis will be rejected at level of significance $\gamma$ if the observed value of $W_{n,\alpha}>\chi^{2}_{\gamma;r}$, the $(1-\gamma)$-th quantile of $\chi^{2}_{r}$ distribution.
Remark 3.1.
The above Wald-type tests coincide with the classical Wald-test at $\alpha=0$, due to the fact that the MDPDE coincides with the MLE at $\alpha=0$.
By the general theory discussed in Basu et al. [5], the above Wald-type test can be shown to be consistent at any fixed alternative for every values of $\alpha\geq 0$. Now, let us consider the contiguous alternative hypotheses of the form
$H_{1,n}:\bm{L\beta}=\bm{l}_{n},\mbox{ where }\bm{l}_{n}=\bm{l}_{0}+n^{-1/2}\bm{d},$ (24)
with $\bm{d}\in\mathbb{R}^{r}-\{\bm{0}_{r}\}$ be a fixed vector. Then, by Theorem 8 of Basu et al. [5], under $H_{1,n}$, $W_{n,\alpha}\xrightarrow{\mathcal{D}}\chi^{2}_{r}(\delta)$, the non-central chi-square distribution with degrees of freedom $r$ and the non-centrality parameter $\delta$, where
$\delta=\frac{1}{nv_{1\alpha}\sigma_{0}^{2}}\bm{d^{T}\left[L\left(\dot{\mu}(\beta_{0})^{T}\dot{\mu}(\beta_{0})\right)^{-1}L^{T}\right]^{-1}d},$ (25)
with $\left(\bm{\beta_{0}}^{T},\sigma_{0}^{2}\right)^{T}\in\bm{\Theta_{0}}$. Hence, an approximate expression of asymptotic contiguous power under $H_{1,n}$ in (24) is given by
$\pi_{n,\alpha}=1-G_{\chi^{2}_{r}(\delta)}\left(\chi^{2}_{\gamma;r}\right),$
where $G_{\chi^{2}_{r}(\delta)}(\cdot)$ denotes the distribution function of $\chi^{2}_{r}(\delta)$.
Let us now consider the setup of Section 2.3 in order to study the robustness of the proposed Wald-type test statistics. We first define the statistical functional corresponding to the test statistics $W_{n,\alpha}$ as given by
$W_{\alpha}(\bm{G})=\left(\bm{LT^{\beta}}_{\alpha}(\bm{G})-\bm{l_{0}}\right)^{T}\left[\bm{L}\left(\frac{\bm{\dot{\mu}}\left(\bm{T^{\beta}}_{\alpha}(\bm{G})\right)^{T}\bm{\dot{\mu}}\left(\bm{T^{\beta}}_{\alpha}(\bm{G})\right)}{nv_{1\alpha}T_{\alpha}^{\sigma^{2}}(\bm{G})}\right)^{-1}\bm{L}^{T}\right]^{-1}\left(\bm{LT^{\beta}}_{\alpha}(\bm{G})-\bm{l_{0}}\right),$
where $\bm{T^{\beta}}_{\alpha}$ and $T_{\alpha}^{\sigma^{2}}$ are the minimum DPD functional corresponding to parameter $\bm{\beta}$ and $\sigma^{2}$, respectively, as defined in Section 2.3.
Let $\bm{\theta_{0}}=\left(\bm{\beta_{0}}^{T},\sigma_{0}^{2}\right)^{T}\in\bm{\Theta}_{0}$ and $\bm{F_{\theta_{0}}}$ be the vector of the model distributions under the null hypotheses. It follows from the theory of general Wald-type tests (see, e.g., [7, 5]) that the first order influence function of the test functional $W_{\alpha}$ is zero when evaluated at $\bm{G}=\bm{F_{\theta_{0}}}$. The corresponding second order IF of the Wald-type test statistic $W_{\alpha}$ at $\bm{G}=\bm{F_{\theta_{0}}}$, when there is contamination only in the $i_{0}$-th observation, is obtained as
$IF_{i_{0}}^{(2)}\left(t_{i_{0}};W_{\alpha},\bm{F_{\theta_{0}}}\right)\\ =\frac{2}{nv_{1\alpha}\sigma_{0}^{2}}IF_{i_{0}}\left(t_{i_{0}};\bm{T^{\beta}}_{\alpha},\bm{F_{\theta_{0}}}\right)^{T}\bm{L^{T}\left[L\left(\dot{\mu}(\beta_{0})^{T}\dot{\mu}(\beta_{0})\right)^{-1}L^{T}\right]^{-1}L}IF_{i_{0}}\left(t_{i_{0}};\bm{T^{\beta}}_{\alpha},\bm{F_{\theta_{0}}}\right).$
Note that, this second order IF directly depends on the IF of the MDPDE functional $\bm{T^{\beta}}_{\alpha}$ of $\bm{\beta}$. Since $IF_{i_{0}}(t_{i_{0}};\bm{T^{\beta}}_{\alpha},\bm{F_{\theta_{0}}})$ is bounded for all $\alpha>0$ and unbounded at $\alpha=0$, so is $IF_{i_{0}}^{(2)}\left(t_{i_{0}};W_{\alpha},\bm{F_{\theta_{0}}}\right)$. This ensures the robustness of the Wald-type test for any $\alpha>0$ as well.
We can also study the robustness of the level and power of these Wald-type tests by examining its level influence function (LIF) and power influence function (PIF). From Theorem 12 of Basu et al. [5], we can conclude that the LIF of the MDPDE based Wald-type test is identically zero whenever the IF of the MDPDE used in constructing the test is bounded, i.e., for all $\alpha>0$. Further, its PIF at level of significance $\gamma$ is given by
$PIF\left(\bm{t};W_{\alpha},\bm{F_{\theta_{0}}}\right)=K_{r}^{*}(\delta)\frac{1}{nv_{1\alpha}\sigma_{0}^{2}}\bm{d^{T}\left[L\left(\dot{\mu}(\beta_{0})^{T}\dot{\mu}(\beta_{0})\right)^{-1}L^{T}\right]^{-1}}IF\left(\bm{t;T_{\alpha},F_{\theta_{0}}}\right),$
where $\delta$ is as given in (25) and
$K_{r}^{*}(s)=e^{-\frac{s}{2}}\sum_{v=0}^{\infty}\frac{s^{v-1}}{v!2^{v}}(2v-s)P\left(\chi^{2}_{r+2v}>\chi^{2}_{\gamma;p}\right).$
Note that, the PIF is linear in the IF of the MDPDE, and hence is bounded for any $\alpha>0$. This implies the robustness of the power of our Wald-type tests.
3.2 Tests for scalar parameter against one sided alternative
Let us now consider the problem of testing a single parameter component, against a one-sided alternative, i.e., testing the hypothesis
$H_{0}:\beta_{k}=\beta_{k}^{0}\mbox{ against }H_{1}:\beta_{k}>\beta_{k}^{0},$ (26)
where $\beta_{k}$ is the $k$-th component of regression parameter $\bm{\beta}$ in our NLR model (3). Under the both sided alternative $H_{1}^{*}:\beta_{k}\neq\beta_{k}^{0}$, the general MDPDE based Wald-type test statistics can be obtained from discussion of the previous subsection, which asymptotically follow $\chi^{2}_{1}$. However, for testing the one-sided alternative in (26), since we are dealing with a scalar parameter, we can use the test statistics
$\widetilde{W}_{n,\alpha}=\frac{\widehat{\beta}_{k\alpha}-\beta_{k}^{0}}{\sqrt{v_{1\alpha}\widehat{\sigma}_{\alpha}^{2}s_{k}}},$ (27)
where $\widehat{\beta}_{k\alpha}$ is the $k$-th component of $\bm{\widehat{\beta}_{\alpha}}$, and $s_{k}$ is the $(k,k)$-th element of the matrix $\bm{\left(\dot{\mu}\left(\widehat{\beta}_{\alpha}\right)^{T}\dot{\mu}\left(\widehat{\beta}_{\alpha}\right)\right)}^{-1}$. It is evident from Theorem 2.1 that, under $H_{0}$ in (26), $\widetilde{W}_{n,\alpha}$ asymptotically follows a standard normal distribution. So, this null hypothesis will be rejected against the one sided alternative $H_{1}$ in (26), at level of significance $\gamma$, if the observed value of $W_{n,\alpha}$ exceeds $z_{\gamma}$, the $(1-\gamma)$-th quantile of $\mathcal{N}(0,1)$.
Further, with respect to the one-sided contiguous alternative hypotheses of the form
$H_{1,n}:\beta_{k}=\beta_{k}^{0}+n^{-1/2}d,\ d>0,$
the large sample approximate expression of the power is given by
$\widetilde{\pi}_{n,\alpha}=P_{H_{1,n}}\left[\widetilde{W}_{n,\alpha}>z_{\gamma}\right]\approx 1-\Phi\left(z_{\gamma}-\sqrt{\frac{1}{\sigma^{2}}\left(1+\frac{\alpha^{2}}{1+2\alpha}\right)^{-3/2}}\frac{d}{\sqrt{ns_{k}}}\right).$
Note that, this asymptotic contiguous power $\widetilde{\pi}_{n,\alpha}$ depends on the covariates and the model mean function only through the quantity $d^{*}=\frac{d}{\sqrt{ns_{k}}}$. Table 2 shows the empirical values of $\widetilde{\pi}_{n,\alpha}$ at 5% level of significance, over different values of tuning parameter $\alpha$ and for different values of $d^{*}$, with $\sigma^{2}=1$. Clearly, for fixed $\alpha$, the power increases with $d^{*}$; in fact $d^{*}=0$ gives the level of the test. But, for fixed $d^{*}>0$, the power of our proposed test decreases with increasing values of $\alpha$, which is quite natural, following the pattern of the ARE of the MDPDE used to construct these tests. Additionally, the loss in power is not quite significant compared to the classical MLE based test (at $\alpha=0$) for small positive values of $\alpha$ (and also for larger $\alpha$ values when $d^{*}$ is large).
Table 2: Asymptotic contiguous power of the Wald-type test for testing $H_{0}$ in (26) against $H_{1,n}$
$d^{*}$ $\alpha$
0 0.1 0.3 0.5 0.7 1
0 0.050 0.050 0.050 0.050 0.050 0.050
1 0.260 0.258 0.247 0.233 0.219 0.201
2 0.639 0.634 0.608 0.574 0.538 0.487
3 0.912 0.909 0.891 0.865 0.833 0.780
5 1.000 1.000 0.999 0.998 0.997 0.991
7 1.000 1.000 1.000 1.000 1.000 1.000
4 Application to the Michaelis-Menten (MM) Model
Let us now apply the proposed robust inference procedure to the popular MM model, as defined by Equation (2) in Section 1. To be consistent with our notation so far, let us denote the observed reaction velocities as $y_{i}$ and the substrate concentrations as $x_{i}$ for the $i$-th observation, and rewrite the MM model equation as
$y_{i}=\mu(x_{i},\bm{\beta})+\epsilon_{i},\ i=1,2,\ldots,n,$ (28)
where $\mu(x_{i},\bm{\beta})=\frac{\beta_{1}x_{i}}{\beta_{2}+x_{i}}$ with $\bm{\beta}=(\beta_{1},\beta_{2})^{T}=(V_{max},K_{m})^{T}$ being the regression parameter vector and the random errors $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, independently for all $i=1,\ldots,n$. As the mean function here is not defined at $\beta_{2}=-x_{i}$, for any $i$, so we must take the parameter space for $\beta_{2}$ to be a subset of $\mathbb{R}\setminus\{-x_{1},\ldots,-x_{n}\}$; in applications, we generally have $\beta_{2}>0$, which ensures that the corresponding parameter space ($\mathbb{R}^{+}$) is independent of the values of the covariate, and hence the resulting mean function always remains well-defined. The parameter space for $\beta_{1}$ may be taken as $\mathbb{R}$.
For the MM model (28), let us denote $\widehat{\bm{\theta}}_{\alpha}=\left(\widehat{\bm{\beta}}_{\alpha}^{T},\widehat{\sigma}_{\alpha}^{2}\right)^{T}=\left(\widehat{\beta_{1}}_{\alpha},\widehat{\beta_{2}}_{\alpha},\widehat{\sigma}_{\alpha}^{2}\right)^{T}$ to be the MDPDE of $\bm{\theta}=\left(\bm{\beta}^{T},\sigma^{2}\right)^{T}$ corresponding to a tuning parameter $\alpha\geq 0$, defined as the minimizer of the objective function in (5) with the particular model specific form of $\mu(x_{i},\bm{\beta})=\frac{\beta_{1}x_{i}}{\beta_{2}+x_{i}}$. The corresponding estimating equations are given by (6) and (7), where we now have $\nabla_{\bm{\beta}}\mu(\bm{x_{i}},\bm{\beta})=\left(\frac{x_{i}}{\beta_{2}+x_{i}},\ -\frac{\beta_{1}x_{i}}{(\beta_{2}+x_{i})^{2}}\right)^{T}$.
4.1 Asymptotic properties
In order to find the asymptotic distributions of the MDPDEs of the parameters of the MM model, we follow the notation of Section 2.2 to get $\bm{\mu(\beta)}=\left(\frac{\beta_{1}x_{1}}{\beta_{2}+x_{1}},\ldots,\frac{\beta_{1}x_{n}}{\beta_{2}+x_{n}}\right)^{T}$ and hence
$\bm{\dot{\mu}(\beta)}=\begin{bmatrix}\frac{x_{1}}{\beta_{2}+x_{1}}&-\frac{\beta_{1}x_{1}}{(\beta_{2}+x_{1})^{2}}\\ \vdots&\vdots\\ \frac{x_{n}}{\beta_{2}+x_{n}}&-\frac{\beta_{1}x_{n}}{(\beta_{2}+x_{n})^{2}}\end{bmatrix}.$ (29)
Then, under the general Conditions $(R1)$ and $(R2)$, it follows from Theorem 2.1 that
$\widehat{\bm{\beta}}_{\alpha}-\bm{\beta}\xrightarrow{\mathcal{D}}\mathcal{N}_{2}\left(\bm{0},v_{1\alpha}\sigma^{2}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)^{-1}\right),$
where
$\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}=\begin{bmatrix}\sum_{i=1}^{n}a_{i}^{2}&-\sum_{i=1}^{n}a_{i}b_{i}\\ \\ -\sum_{i=1}^{n}a_{i}b_{i}&\sum_{i=1}^{n}b_{i}^{2}\end{bmatrix},$ (30)
with $a_{i}=\frac{x_{i}}{\beta_{2}+x_{i}}$, $b_{i}=\frac{\beta_{1}x_{i}}{(\beta_{2}+x_{i})^{2}}$. Upon simplifications, we get
$\widehat{\beta_{1}}_{\alpha}-\beta_{1}\xrightarrow{\mathcal{D}}\mathcal{N}\left(0,\sigma_{\beta_{1}}^{2}\right)\mbox{ and }\widehat{\beta_{2}}_{\alpha}-\beta_{2}\xrightarrow{\mathcal{D}}\mathcal{N}\left(0,\sigma_{\beta_{2}}^{2}\right),$
where $\sigma_{\beta_{1}}^{2}=\frac{v_{1\alpha}\sigma^{2}\sum_{i=1}^{n}b_{i}^{2}}{\sum_{i=1}^{n}a_{i}^{2}\sum_{i=1}^{n}b_{i}^{2}-\left(\sum_{i=1}^{n}a_{i}b_{i}\right)^{2}}$ and $\sigma_{\beta_{2}}^{2}=\frac{v_{1\alpha}\sigma^{2}\sum_{i=1}^{n}a_{i}^{2}}{\sum_{i=1}^{n}a_{i}^{2}\sum_{i=1}^{n}b_{i}^{2}-\left(\sum_{i=1}^{n}a_{i}b_{i}\right)^{2}}$.
In order to help verification of the required Conditions (R1)-(R2) for a given MM model, we have proved the following propositions; see Appendices B.2 and B.3 for their proofs.
Proposition 1.
Under the MM model given in (28), if the parameter space for $\beta_{2}$ is $[0,\infty)$ and $x_{i}>0$ for all $i$, then Condition (R1) is satisfied.
Proposition 2.
In order to satisfy Condition (14) in (R2) for the MM model given by (28), it is necessary that the values $x_{i}$’s are not all same, $n\geq 2$ and $\beta_{1}\neq 0$.
Note that, Proposition 1 provides a set of easily verifiable sufficient conditions for (R1) which are relatively stricter and not necessary for (R1). It may be possible to verify (R1) directly even if these sufficient conditions are not satisfied (although they are assumed to be true in most applications of the MM model). On the other hand, Proposition 2 provides a set of conditions which are necessary for (R2) to hold.
4.2 Influence function
We will now study the IFs of the MDPDEs of the parameters of the MM model following the general theory developed in Section 2.3. Using (17) and (30), the expressions of IFs of minimum DPD functionals $T_{\alpha}^{\beta_{1}}$ and $T_{\alpha}^{\beta_{2}}$ corresponding to the parameters $\beta_{1}$ and $\beta_{2}$ of the MM model (28), respectively, simplify to
$IF_{i_{0}}(t_{i_{0}},T_{\alpha}^{\beta_{1}},\bm{F_{\theta}})=\\ \frac{(1+\alpha)^{\frac{3}{2}}(t_{i_{0}}-\mu(x_{i_{0}},\bm{\beta}))e^{-\frac{\alpha}{2\sigma^{2}}(t_{i_{0}}-\mu(x_{i_{0}},\bm{\beta}))^{2}}}{\det\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)}\frac{\beta_{1}^{2}x_{i_{0}}}{(\beta_{2}+x_{i_{0}})^{2}}\sum_{j=1}^{n}\frac{x_{j}^{2}(x_{i_{0}}-x_{j})}{(\beta_{2}+x_{j})^{4}},$ (31)
$IF_{i_{0}}(t_{i_{0}},T_{\alpha}^{\beta_{2}},\bm{F_{\theta}})=\\ \frac{(1+\alpha)^{\frac{3}{2}}(t_{i_{0}}-\mu(x_{i_{0}},\bm{\beta}))e^{-\frac{\alpha}{2\sigma^{2}}(t_{i_{0}}-\mu(x_{i_{0}},\bm{\beta}))^{2}}}{\det\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)}\frac{\beta_{1}x_{i_{0}}}{(\beta_{2}+x_{i_{0}})^{2}}\sum_{j=1}^{n}\frac{x_{j}^{2}(x_{i_{0}}-x_{j})}{(\beta_{2}+x_{j})^{3}},$ (32)
when there is contamination only in $i_{0}$-th observation at the contamination point $t_{i_{0}}$. For infinitesimal contamination in all observations at the contamination points in $\bm{t}=(t_{1},\cdots,t_{n})^{T}$, these IFs are given by
$IF(\bm{t},T_{\alpha}^{\beta_{k}},\bm{F_{\theta}})=\sum_{i=1}^{n}IF_{i}(t_{i},T_{\alpha}^{\beta_{k}},\bm{F_{\theta}}),\ \ k=1,2.$ (33)
For illustrations, we have plotted these IFs of the MDPDEs under the MM-model (28), by generating data from the following two particular setups:
  • Setup 1: $n=50,\ (\beta_{1},\beta_{2},\sigma)=(5,1,1),$ and $x_{i}=i$, for all $i=1,2,\ldots,n$.
  • Setup 2: $n=50,\ (\beta_{1},\beta_{2},\sigma)=(50,2,2),$ and $(x_{1},x_{2},\ldots,x_{n})$ is a random sample of size $n$ from Uniform$(0,40)$.
The resulting plots of the IFs for contaminations in all observations are presented in Figure 1 for both Setups 1-2, respectively; those obtained for contamination in particular observation are provided in Appendix C.1. To plot the IFs in (33) (in Figure 1) on a 2-D surface, we have considered same contamination point for all observations, i.e., the plots are obtained for the IFs at contamination points in $\bm{t}^{*}=(t,\ldots,t)$. From all these figures, it is clear that the IFs of the MDPDEs under the MM model are bounded for $\alpha>0$ and are redecending in nature for increasing values of $\alpha$. For $\alpha=0$, the IFs are unbounded as expected, since the MDPDE with $\alpha=0$ is the same as the non-robust MLE also for the MM models.
(a) $IF(\bm{t}^{*},T_{\alpha}^{\beta_{1}},\bm{F_{\theta}})$, under Setup 1
(b) $IF(\bm{t}^{*},T_{\alpha}^{\beta_{2}},\bm{F_{\theta}})$, under Setup 1
(c) $IF(\bm{t}^{*},T_{\alpha}^{\beta_{1}},\bm{F_{\theta}})$, under Setup 2
(d) $IF(\bm{t}^{*},T_{\alpha}^{\beta_{2}},\bm{F_{\theta}})$, under Setup 2
Figure 1: IFs of the MDPDEs under the MM model, with contamination in all observations, at the same contamination point $t$ [i.e., $\bm{t}^{*}=(t,\ldots,t)$]
4.3 Finite sample performances of the MDPDE
We now present an extensive simulation study to justify the improved performances of the MDPDEs of parameters of the MM model (28). For this purpose we have simulated data following the two setups (Setup 1 and Setup 2) as mentioned in the previous Section 4.2. Additionally, we have incorporated data contaminations by multiplying the observed values of response or covariate (or both) by a fixed number, say $C$, in a certain proportion (say $e_{c}$) of cases selected randomly from the full sample. We report the results for the choice of $C=5$ under Setup 1 and for $C=2$ under Setup 2.
For each setup, we present the empirical absolute bias (EBias) and the empirical mean square error (EMSE) of the MDPDE and different other robust estimators (as described in Appendix A.2) of the parameters of the MM model (28), based on $1000$ replications for pure data as well as contaminated data with different choices of the contamination proportion $e_{c}$. The results for $e_{c}=10\%$ (light contamination) and $e_{c}=40\%$ (heavy contamination) are reported in Tables 3-4, along with the same obtained for the pure data ($e_{c}=0\%$); the results obtained for $e_{c}=20\%$ and $30\%$ (moderate contamination) are reported in Appendix C.2 for brevity. In particular, we have compared the performances of the MDPDEs at different values of $\alpha$ with the OLS (which is the same as the MDPDE at $\alpha=0$), the MOM estimator, the KPS estimator with different choice of its tuning parameter $w$, and M-estimators based on the Huber’s and the Tukey’s loss/weight functions.
In addition, we have also computed the average prediction errors for the model fitted with different estimators which we define based only on the clean observations as given by
$\frac{1}{1000}\sum_{r=1}^{1000}\frac{1}{|\mathcal{A}_{r}|}\sum_{i\in\mathcal{A}_{r}}\left(y_{i}-\mu\left(\widehat{\bm{\beta}}^{(r)},x_{i}\right)\right)^{2},$
[…]
$\left.r\mbox{-th replication,}\ i=1,2,\ldots,n\right\}$ and $\widehat{\bm{\beta}}^{(r)}$ be parameter estimate obtained at $r$-th replication.
Based on all these numerical results, we can easily see that the MDPDEs with suitable values of $\alpha$ perform the best among all the competitors considering all the cases of simulation setups. For lower contamination proportions, a close competitor of our MDPDE is the KPS estimator and the Tukey’s M estimator, but their performance deteriorates as contamination proportion increases. This deterioration is more significant when there are contamination in both response and covariate. Further, from Tables 5 and 10 it is clear that, for increasing proportion of contamination, the average prediction error gets minimized for a larger value of the MDPDE tuning parameter $\alpha$.
Table 3: Empirical absolute bias and empirical MSE obtained by different estimators in simulations with Setup 1 and different amounts of contamination (the minimum values obtained for each parameter are highlighted in bold font)
Outlier Direction No outlier (pure data) Contamination only in response Contamination in both response and covariate
Outlier proportion 10% 40% 10% 40%
$\beta_{1}$ $\beta_{2}$ $\sigma$ $\beta_{1}$ $\beta_{2}$ $\sigma$ $\beta_{1}$ $\beta_{2}$ $\sigma$ $\beta_{1}$ $\beta_{2}$ $\sigma$ $\beta_{1}$ $\beta_{2}$ $\sigma$
Empirical Bias
OLS 0.0110 0.0677 2.6532 2.3313 9.1277 2.1085 69.35 377.97 36.35 119.67
MOM 0.0814 0.1973 85.37 1.1631 59.26 6.7111 0.4602 205.45 7.1474 363.16
KPS(w=0.5) 0.0143 0.0728 0.0083 0.0466 0.3470 0.7564 0.0184 0.0734 31.80 116.04
KPS(w=1) 0.0256 0.1087 0.0200 0.0964 0.0300 0.1732 0.0220 0.1006 0.0789 0.1252
KPS(w=1.5) 0.0533 0.2256 0.0394 0.1717 0.0036 0.0036 0.0410 0.1767 0.0276 0.0815
M-Est (Huber) 0.0141 0.0809 0.2059 0.0326 4.4738 1.2853 0.4880 1.0417 11.23 9.0000
M-Est (Tukey) 0.0135 0.0800 0.0104 0.0682 2.5193 1.4471 0.0137 0.0778 11.25 8.9182
MDPDE($\alpha$=0.05) 0.0110 0.0677 0.0232 0.4978 0.0948 1.3548 7.7955 0.4942 8.6529 12.79 63.72 1.0553 31.68 96.26 4.9905
MDPDE($\alpha$=0.1) 0.0112 0.0690 0.0224 0.0010 0.0271 0.0037 2.3448 1.1511 4.0321 0.0173 0.0721 0.0003 15.69 51.07 2.3681
MDPDE($\alpha$=0.2) 0.0120 0.0723 0.0223 0.0056 0.0538 0.0106 0.0156 0.0208 0.0957 0.0124 0.0733 0.0098 0.2449 0.8665 0.1055
MDPDE($\alpha$=0.3) 0.0130 0.0766 0.0238 0.0098 0.0704 0.0064 0.0063 0.0496 0.1081 0.0128 0.0782 0.0058 0.0634 0.2640 0.1095
MDPDE($\alpha$=0.5) 0.0156 0.0861 0.0299 0.0143 0.0883 0.0027 0.0122 0.0851 0.2053 0.0153 0.0897 0.0032 0.0267 0.1263 0.2087
MDPDE($\alpha$=0.7) 0.0195 0.0986 0.0384 0.0179 0.1006 0.0106 0.0173 0.1032 0.3147 0.0189 0.1019 0.0109 0.0274 0.1312 0.3180
MDPDE($\alpha$=1) 0.0272 0.1259 0.0514 0.0243 0.1226 0.0198 0.0210 0.1174 0.4614 0.0248 0.1226 0.0200 0.0287 0.1377 0.4646
Empirical MSE
OLS 0.0422 0.2796 15.91 143.76 130.23 221.72 5931 180529 1358 15184
MOM 0.1283 0.9449 7.1E+06 33.18 2.2E+06 533.73 0.9885 50252 63.87 1.2E+7
KPS(w=0.5) 0.0483 0.3253 0.0575 0.4145 1.0585 2.6777 0.0547 0.3925 1376 18957
KPS(w=1) 0.0677 0.4583 0.0691 0.4915 0.1233 0.9721 0.0681 0.4817 0.1185 0.9277
KPS(w=1.5) 0.1121 0.9119 0.0977 0.7668 0.1095 0.8442 0.0974 0.7585 0.1028 0.8029
M-Est (Huber) 0.0446 0.3071 0.1032 0.4390 31.11 10.15 0.4239 3.7898 127.64 81.00
M-Est (Tukey) 0.0447 0.3084 0.0504 0.3715 27.40 11.60 0.0494 0.3587 129.72 80.25
MDPDE($\alpha$=0.05) 0.0423 0.2808 0.0100 0.9278 1.9099 5.0569 94.07 160.72 77.42 473.90 11987 2.8641 1041 9818 25.92
MDPDE($\alpha$=0.1) 0.0426 0.2842 0.0101 0.0538 0.4031 0.0144 59.55 463.88 38.60 0.0492 0.3540 0.0150 532.03 5752 12.40
MDPDE($\alpha$=0.2) 0.0438 0.2958 0.0107 0.0525 0.3904 0.0124 0.1546 0.7214 0.3195 0.0499 0.3605 0.0126 6.6965 84.18 0.2012
MDPDE($\alpha$=0.3) 0.0457 0.3142 0.0116 0.0533 0.3928 0.0132 0.0953 0.7249 0.0961 0.0518 0.3771 0.0133 1.3977 21.98 0.0552
MDPDE($\alpha$=0.5) 0.0513 0.3641 0.0140 0.0576 0.4302 0.0158 0.0906 0.7270 0.0739 0.0569 0.4219 0.0159 0.0875 0.6895 0.0757
MDPDE($\alpha$=0.7) 0.0577 0.4189 0.0169 0.0634 0.4804 0.0189 0.0922 0.7410 0.1373 0.0627 0.4717 0.0189 0.0901 0.7245 0.1397
MDPDE($\alpha$=1) 0.0695 0.5153 0.0214 0.0730 0.5654 0.0230 0.0946 0.7683 0.2605 0.0726 0.5597 0.0230 0.0934 0.7629 0.2637
Table 4: Empirical absolute bias and empirical MSE obtained by different estimators in simulations with Setup 2 and different amounts of contamination (the minimum values obtained for each parameter are highlighted in bold font)
Outlier Direction No outlier (pure data) Contamination only in response Contamination in both response and covariate
Outlier proportion 10% 40% 10% 40%
$\beta_{1}$ $\beta_{2}$ $\sigma$ $\beta_{1}$ $\beta_{2}$ $\sigma$ $\beta_{1}$ $\beta_{2}$ $\sigma$ $\beta_{1}$ $\beta_{2}$ $\sigma$ $\beta_{1}$ $\beta_{2}$ $\sigma$
Empirical Bias
OLS 0.0053 0.0021 5.3409 0.1399 20.73 0.2362 11.99 2.7578 48.00 11.62
MOM 0.0227 0.0089 1.3296 0.0845 22.67 0.7303 2.8428 0.6894 55.55 12.25
KPS(w=0.5) 0.0154 0.0047 0.0061 0.0055 0.5786 0.5918 0.0221 0.0128 0.3363 0.1575
KPS(w=1) 0.0359 0.0096 0.0225 0.0042 0.2046 0.0864 0.0114 0.0001 0.2167 0.1067
KPS(w=1.5) 0.0637 0.0176 0.0455 0.0107 0.0415 0.0197 0.0404 0.0092 0.1108 0.0518
M-Est (Huber) 0.0403 0.0147 0.3804 0.0233 8.4130 0.1436 0.8063 0.1190 37.64 6.9652
M-Est (Tukey) 0.0423 0.01525 0.0364 0.0103 4.7816 0.0202 0.0197 0.0031 37.61 7.0127
MDPDE($\alpha$=0.05) 0.0056 0.0023 0.0521 0.0155 0.0697 0.6136 2.5806 2.1279 19.16 0.4343 0.0112 0.9795 24.08 4.5723 19.37
MDPDE($\alpha$=0.1) 0.0059 0.0025 0.0515 0.0168 0.0089 0.0212 0.4501 0.4815 4.0300 0.0430 0.0218 0.0012 5.8616 0.7556 6.4756
MDPDE($\alpha$=0.2) 0.0074 0.0030 0.0534 0.0007 0.0017 0.0272 0.0469 0.0288 0.1562 0.0167 0.0095 0.0192 0.1414 0.0778 0.2317
MDPDE($\alpha$=0.3) 0.0094 0.0036 0.0583 0.0047 0.0002 0.0205 0.0240 0.0144 0.2090 0.0084 0.0059 0.0158 0.1110 0.0537 0.2486
MDPDE($\alpha$=0.5) 0.0141 0.0048 0.0730 0.0089 0.0012 0.0049 0.0044 0.0063 0.4105 0.0004 0.0028 0.0026 0.0706 0.0355 0.4303
MDPDE($\alpha$=0.7) 0.0185 0.0058 0.0914 0.0126 0.0023 0.0097 0.0004 0.0044 0.6267 0.0056 0.0007 0.0107 0.0565 0.0292 0.6363
MDPDE($\alpha$=1) 0.0269 0.0081 0.1197 0.0206 0.0051 0.0257 0.0035 0.0031 0.9167 0.0159 0.0030 0.0257 0.0454 0.0242 0.9152
Empirical MSE
OLS 0.2960 0.0243 36.82 0.6774 454.68 1.3956 185.17 14.72 2515.37 180.11
MOM 0.6553 0.0534 6.1540 0.2222 648.93 7.1888 27.66 2.6532 3510.54 248.81
KPS(w=0.5) 0.3349 0.0271 0.3595 0.0290 2.0150 0.7564 0.3629 0.0298 17.01 1.6779
KPS(w=1) 0.4660 0.0373 0.4538 0.0367 0.8494 0.0884 0.4522 0.0368 0.7110 0.0748
KPS(w=1.5) 0.7114 0.0554 0.6407 0.0522 0.6520 0.0516 0.6290 0.0512 0.6813 0.0600
M-Est (Huber) 0.3165 0.0267 0.5872 0.0404 146.57 1.6212 1.1252 0.0586 1468.90 52.26
M-Est (Tukey) 0.3173 0.0267 0.3406 0.0284 159.54 1.8608 0.3476 0.0293 1478.86 53.08
MDPDE($\alpha$=0.05) 0.2963 0.0245 0.0445 0.5644 0.06908 4.9107 416.30 42.78 611.86 2.8726 0.0941 7.3089 1537.79 127.73 570.47
MDPDE($\alpha$=0.1) 0.2986 0.0248 0.0450 0.3367 0.02840 0.0548 56.44 4.6886 124.26 0.3459 0.0304 0.0603 488.85 85.12 184.64
MDPDE($\alpha$=0.2) 0.3071 0.0257 0.0474 0.3401 0.02841 0.0517 0.6918 0.0553 0.4673 0.3468 0.0293 0.0529 2.0120 0.1241 0.4688
MDPDE($\alpha$=0.3) 0.3203 0.0270 0.0512 0.3530 0.02962 0.0550 0.5915 0.0506 0.1591 0.3574 0.0302 0.0555 0.6441 0.0628 0.1901
MDPDE($\alpha$=0.5) 0.3561 0.0304 0.0615 0.3873 0.03285 0.0645 0.5914 0.0487 0.3042 0.3891 0.0332 0.0649 0.6338 0.0577 0.3275
MDPDE($\alpha$=0.7) 0.4013 0.0345 0.0738 0.4273 0.03652 0.0746 0.6063 0.0499 0.5570 0.4279 0.0368 0.0748 0.6431 0.0574 0.5724
MDPDE($\alpha$=1) 0.4777 0.0411 0.0920 0.4927 0.04275 0.0881 0.6240 0.0513 1.0455 0.4916 0.0429 0.0881 0.6542 0.0571 1.0430
Table 5: Average prediction errors obtained by different estimators and different amounts of contamination (the minimum values obtained for each parameter are highlighted in bold font)
Setup Setup 1 Setup 2
Outlier direction none response response + covariate none response response + covariate
Outlier proportion 0% 10% 40% 10% 40% 0% 10% 40% 10% 40%
OLS 0.961 4.865 56.80 6.068 16.40 3.828 23.91 296.80 32.98 245.91
MOM 1.974 14.02 4504.9 2.922 18.50 4.033 6.618 341.71 10.20 388.62
KPS(w=0.5) 0.967 1.017 551.94 0.963 3.874 3.845 3.829 9.372 3.831 5.760
KPS(w=1) 0.983 0.997 2.155 0.978 3.817 3.915 3.876 3.872 3.878 3.852
KPS(w=1.5) 1.009 1.528 5.944 1.939 1.554 4.032 3.976 3.775 3.967 3.806
M-Est (Huber) 0.963 1.003 15.67 1.048 42.94 3.838 3.998 78.60 4.055 244.95
M-Est (Tukey) 0.963 0.958 10.61 0.957 43.92 3.838 3.818 64.45 3.821 243.90
MDPDE($\alpha$=0.05) 0.961 3.147 54.59 6.333 18.58 3.829 16.13 315.87 21.39 242.94
MDPDE($\alpha$=0.1) 0.961 0.965 52.27 1.819 16.72 3.830 3.817 366.36 4.638 293.84
MDPDE($\alpha$=0.2) 0.962 0.960 47.68 0.957 13.97 3.835 3.819 146.64 3.819 221.57
MDPDE($\alpha$=0.3) 0.964 0.968 42.11 0.959 11.58 3.842 3.926 20.366 3.825 97.43
MDPDE($\alpha$=0.5) 0.971 0.965 4.665 0.964 4.348 3.861 3.844 3.765 3.842 4.427
MDPDE($\alpha$=0.7) 0.984 1.000 0.943 0.972 1.056 3.886 3.866 3.767 3.864 3.786
MDPDE($\alpha$=1) 0.995 0.987 0.963 0.985 0.943 3.929 3.902 3.776 3.900 3.792
We would like to particularly mention that the median of means (MOM) method fails to perform well for almost all the contamination schemes. So, we have not considered this MOM estimator for further real data applications presented in this paper.
4.4 Wald-type tests for validity of the Michaelis-Menten model
In the context of our MM model (28), we assume $x_{i}>0$, for all $i=1,2,\ldots,n$, as $x_{i}$’s are the substrate concentrations in a reaction. Now, $\beta_{2}=0$ implies that the mean/regression function is simply $\mu(x_{i},\bm{\beta})=\beta_{1}$, a constant, which is independent of the covariate $x_{i}$; Consequently $y_{i}$’s simply become IID observations from a normal distribution with constant mean $\beta_{1}$ and variance $\sigma^{2}$. So, to test the validity of the MM model assumption it is enough to test
$H_{0}:\beta_{2}=0\mbox{ against }H_{1}:\beta_{2}>0.$ (34)
For this test we can use the test statistic $\widetilde{W}_{n,\alpha}$ given in (27), i.e.,
$\widetilde{W}_{n,\alpha}=\frac{\widehat{\beta}_{2\alpha}}{\sqrt{v_{1\alpha}\widehat{\sigma}_{\alpha}^{2}s_{22}}}\xrightarrow{\mathcal{D}}\mathcal{N}(0,1),\mbox{ under }H_{0}.$ (35)
Then, for testing $H_{0}$ against the one sided alternative $H_{1}$ in (34), $H_{0}$ will be rejected at level of significance $\gamma$ if the observed value of $\widetilde{W}_{n,\alpha}$ exceeds $z_{\gamma}$, the $(1-\gamma)$-th quantile of $\mathcal{N}(0,1)$. Similarly, one can construct Wald-type tests for testing of the parameter $\beta_{1}$ against one sided or both sided alternatives.
We have conducted simulation studies to study the performance of the Wald-type test for the above discussed testing problem. In particular, we have computed the empirical level (ELevel) and the empirical power (EPower) of the test for various values of the tuning parameter $\alpha$ based on 1000 replications of simulated data; the simulation schemes for computations of ELevel and EPower are taken as follows:
For ELevel: We have simulated samples as per Setup 2, but with $(\beta_{1},\beta_{2},\sigma)=(20,0,1)$ and different sample sizes ($n$). So, $y_{i}$’s are IID $\mathcal{N}(20,1)$, i.e., the MM model is not significant for these (pure) data. For introducing contamination, $e_{c}$ proportion of $y_{i}$’s are randomly selected and each $y_{i}$ is replaced by a random number between $z_{i}$ and $y_{i}$, where $z_{i}=\frac{20x_{i}}{5+x_{i}}$ is the true mean function of an MM model with parameters $\beta_{1}=20$, $\beta_{2}=5$. We have chosen such a contamination scheme, so as to make the (non-robust) estimates of $\beta_{2}$ likely to be larger than 0, so that a non-robust test will reject the null hypothesis more than expected due to the effect of data contamination, although the null hypothesis is true for the actual generated samples. The ideal values of ELevel, measured as the proportion of rejections, should be close to the level of the test, which is taken to be $5\%$ in our experiments.
For EPower: We have again simulated data following Setup 2 but with $(\beta_{1},\beta_{2},\sigma)=(20,1,2)$ and different sample sizes ($n$). For introducing contamination, $e_{c}$ proportion of $y_{i}$’s are randomly selected and each $y_{i}$ is replaced by a random number generated from Uniform$(0,20)$. In this case the null hypothesis is actually false, i.e., the MM model is indeed valid. So, empirical power, measured as the proportion of rejections, should be high (close or equal to 1) for a good stable test.
The resulting values of the empirical levels and powers are plotted against the sample sizes for various values of tuning parameter $\alpha$ and various proportions ($e_{c}$) of contamination, which are presented in Figures 2 and 3, respectively.
One can observe from Figure 2 that, under pure data the empirical levels of Wald test (at $\alpha=0$) and all the MDPDE based Wald-type tests with $\alpha>0$ are quite close to each other and fluctuate closely around the desired $5\%$ level of significance. Further, for contaminated data, the empirical levels of the MDPDE based Wald-type tests with moderate or large values of $\alpha$ $(\alpha\geq 0.3)$ tend to stabilize near $5\%$ level of significance, as sample size increases. But the empirical levels of the MLE based Wald test fail to converge to the desired level in the presence of contamination due to its non-robust nature.
(a) ELevel, 0% contamination (pure data)
(b) ELevel, 10% contamination
(c) ELevel, 20% contamination
(d) ELevel, 30% contamination
Figure 2: Empirical levels (ELevel) of the robust Wald-type tests
Figure 3 illustrates that, under pure data, the empirical powers of the Wald test and MDPDE based the Wald-type tests are all very close to 1 at a moderate sample size of $n=50$ or $60$, and clearly converge to 1 quite fast as the sample size increases. But for contaminated data, the empirical powers of the Wald tests decrease considerably with increase in the proportion of contamination and the rate of consistency slows down significantly. However, the empirical powers of the MDPDE based Wald-type tests with $\alpha\geq 0.3$ decreases only a little at higher contamination proportions but then also they regain their convergence rate to one very quickly as the sample size increases. These figures clearly illustrate significantly improved stability of our proposed MDPDE based Wald-type tests with increasing values of $\alpha>0$ under possible data contamination.
(a) EPower, 0% contamination (pure data)
(b) EPower, 10% contamination
(c) EPower, 20% contamination
(d) EPower, 30% contamination
Figure 3: Empirical powers (EPower) of the robust Wald-type tests
5 On the choice of optimum tuning parameter
It is well studied that the tuning parameter $\alpha$ of the MDPDE acts as a trade-off between robustness and efficiency, in the sense that the robustness of the estimator, under data contamination, increases at the cost of loss in (asymptotic) efficiency under pure data with increasing values of $\alpha$. The same is also clearly observed for the MDPDEs and the associated Wald-types tests under our NLR model (3) through the extensive numerical illustrations provided in the previous sections. So, while applying these DPD based estimation and testing methods to a real dataset, it is very important to choose an optimal value of $\alpha$ providing the best possible ‘compromise’ between robustness and efficiency depending on the (unknown) amount of contamination in the given dataset.
Under the setup of IID data, Warwick and Jones [39] suggested to use the optimal value of $\alpha$ that minimizes an asymptotic approximation of the MSE of the MDPDEs. Ghosh and Basu [15] extended this idea for the general INH setup, with detailed illustrations for the linear regression model. Following the same idea, an optimal value of $\alpha$ for our NLR model (3) can be obtained by minimizing an estimate of the quantity
$E(\bm{\widehat{\theta}_{\alpha}}-\bm{\theta}^{*})^{T}(\bm{\widehat{\theta}_{\alpha}}-\bm{\theta}^{*})=(\bm{\theta_{\alpha}}-\bm{\theta}^{*})^{T}(\bm{\theta_{\alpha}}-\bm{\theta}^{*})+n^{-1}Trace(\bm{\Psi_{n}^{-1}(\bm{\theta_{\alpha}})\Omega_{n}(\bm{\theta_{\alpha}})\Psi_{n}^{-1}(\bm{\theta_{\alpha}})}),$
where $\bm{\theta}^{*}$ is the target (true) parameter value, $\bm{\theta_{\alpha}}$ is the minimum DPD functional given in (16) and $\bm{\Psi_{n},\Omega_{n}}$ are as given in (11). We may estimate this MSE by replacing $\bm{\theta_{\alpha}}$ by $\bm{\widehat{\theta}_{\alpha}}$ and $\bm{\theta}^{*}$ by a pilot estimator, and then minimize the resulting estimated MSE by grid search over a suitably chosen grid of $\alpha\in[0,1]$. This method crucially depends on the choice of the pilot estimator, although it is shown to work remarkably well for a properly chosen pilot.
In order to remove the dependence on the pilot estimate, Basak et al. [4] proposed an iterative version of the above procedure, namely the Iterated Warwick-Jones (IWJ) method, where they have proposed to use the above procedure iteratively with the pilot estimate being taken as the MDPDE with the tuning parameter obtained in the previous iteration (as optimal). They have numerically illustrated that, starting with the MDPDE with any $\alpha>0$ as the pilot estimate, their IWJ algorithm eventually converges (very fast) to an optimal value of $\alpha$, which is independent of the initial choice of the pilot estimate. We have followed this IWJ approach in order to find the optimal values of $\alpha$ in all our real data examples presented in the following section.
6 Real data examples
Here we present the application of our proposed robust inference methodologies to fit the MM model (28) to some interesting real-life datasets, and compare them with the fits obtained by the existing robust estimators (except for the MOM estimator for the reason discussed in Section 4.3). However, for real data, we are unaware of the exact amount of contaminated observations and hence the prediction accuracy of a fitted model is assessed by the $p\%$ trimmed average prediction error (TAPE), obtained after the largest $p\%$ prediction errors (squared residuals) are being trimmed.
6.1 ONPG data
Our first example is the OPNG dataset developed by Aledo [1]. A batch of second year biochemistry students of the University of Málaga, Spain, using $\beta$-galatosidase (EC. 3.2.1.23) as an enzyme model, carried out different experiments on the hydrolysis of o-nitrophenyl-$\beta$-D-galactopyranoside (ONPG) to galactose and o-nitrophenol, which is an enzyme-catalyzed reaction. Our data, obtained from the R package renz [1], consist of observations from 10 such reactions with different ONPG concentrations where the initial rates (velocity) of these 10 reactions were recorded separately by 8 groups of students. Here, for our illustration purpose, we consider the data obtained by the second group of students, which contains some outlying observations (among the total of $n=10$ observations).
Since these observations are obtained from enzyme-catalyzed reactions, we have fitted the MM model (28) to these data with the initial reaction rates (velocities) as the response and the ONPG concentration as the covariate, using our proposed MDPDEs as well as its robust competitors. The parameter estimates and their (estimated) asymptotic standard errors are given in Table 6 along with the $30\%$ TAPE based on the fitted models for each method. The resulting fitted lines for some selected methods are also shown in Figure 4, along with the scatter plot of the actual observations. For the selection of optimum $\alpha$, we have used the IWJ method as described in Section 5, which yields $\widehat{\alpha}=0.26$, in just one or two iterations for any choice of pilot $\alpha>0$.
Table 6: Parameter estimates of the MM model fitted to the ONPG data, along with their asymptotic standard errors (in parenthesis) and 30% TAPE. [For MDPDE, $\widehat{\alpha}=0.26$ is the optimum choice obtained by IWJ method]
MDPDE
$\alpha$ 0 0.05 0.1 0.2 0.26 0.5 0.7 1
$\beta_{1}$ 165.004 139.908 134.780 134.388 128.515 128.571 128.612 128.675
(55.673) (36.349) (6.736) (6.723) (1.621) (1.901) (2.162) (2.552)
$\beta_{2}$ 12.350 7.726 4.096 4.008 2.582 2.587 2.590 2.596
(9.480) (5.308) (0.653) (0.644) (0.118) (0.138) (0.157) (0.186)
$\sigma$ 21.207 20.687 5.695 5.664 1.667 1.847 1.995 2.180
(201.125) (192.036) (14.687) (14.954) (1.324) (1.786) (2.235) (2.890)
TAPE 85.977 70.492 12.149 11.582 2.005 2.007 2.010 2.014
KPS M-est M-est
w=0.5 w=1 w=1.5 Tukey Huber
$\beta_{1}$ 99.010 101.006 100.861 132.957 136.401
(4.095) (3.330) (4.055) (5.182) (5.888)
$\beta_{2}$ 1.947 1.861 1.850 3.732 4.556
(0.311) (0.239) (0.291) (0.476) (0.609)
TAPE 38.075 47.358 47.650 10.350 14.461
We can see from Table 6 and Figure 4 that the classical OLS fit is heavily affected by the presence of outliers in the data. The KPS method is ignoring the rightmost 2 observations even though they are not outlying observations. Our proposed MDPDE with the optimum value of $\alpha=0.26$ provides much better and stable parameter estimates with lower prediction error for most of the data; in fact, all the values of $\alpha\geq\widehat{\alpha}$ yield similar estimators and model fits. MDPDE with $\alpha=0.1$ and Tukey’s M-estimator also yield reasonable model fits but higher TAPE compared to the MDPDE with $\alpha=0.26$.
Figure 4: ONPG data and the plot of the fitted MM models using different methods
Subsequently, we have also tested for the validity of the Michaelis-Menten model (i.e, test of $\beta_{2}$ following the hypothesis in (34)) for this dataset using the classical Wald test and the proposed MDPDE based Wald-type test. The observed values of the test statistic $\widetilde{W}_{n,\alpha}$, described in (35) in Section 4.4, are given in Table 7 along with the associated (one-sided) p-values.
Table 7: Observed values of test statistic $\widetilde{W}_{n,\alpha}$ for ONPG dataset, and the associated p-values
$\alpha$ 0 (Wald test) 0.05 0.1 0.2 0.26 0.5 0.7 1
$\widetilde{W}_{n,\alpha}$ 1.303 1.456 6.271 6.227 21.954 18.740 16.489 13.982
p-value 0.096 0.073 0.000 0.000 0.000 0.000 0.000 0.000
It is evident that the classical Wald test (at $\alpha=0$) fails to reject the null hypothesis in (34) at $5\%$ level of significance and hence fails to justify the validity of the MM model for this particular dataset, due to the presence of outliers. However, our MDPDE based Wald-type tests with $\alpha\geq 0.1$ provide robust inferences by rejecting the null hypothesis in (34) and hence justifying the validity of the MM model. In fact, these Wald-type tests at the optimum value of $\alpha$ and beyond (i.e., for $\alpha\geq 0.26$) all reject the null hypothesis at any reasonable level of significance, successfully suppressing any effect of outlying observations present in the dataset.
6.2 Drug concentration data
Our second dataset, taken from Tabatabai et al. [36], contains a set of responses for 7 different concentrations of an agonist in a functional assay. Among the 7 observations, the 5-th one is clearly seen to be an outlier; see the scatter plot in Figure 5. Once again we have fitted the MM model (28) for this dataset using the proposed MDPDE, as well as its existing competitors. The parameter estimates are given in the Table 8 along with the asymptotic standard errors and $30\%$ TAPEs for the fitted models, while the model fits are illustrated in Figure 5. The optimum value of $\alpha$ in our MDPDEs, using the IWJ method, turns out to be $\widehat{\alpha}=0.39$ for this dataset.
Like the previous example, here also we can see that the proposed MDPDE at the optimum $\alpha=0.39$ (in fact for all $\alpha\geq 0.39$) yields robust parameter estimates and model fits ignoring the effects of the outlier. Tukey’s M-estimator still remains a close contender with similar parameter estimate with a lower TAPE, but form the Figure 5 it can be clearly observed that it does not provide a fit as good as the MDPDEs with $\alpha\geq 0.39$. Also, the KPS estimators for both $w=0.05$ and $w=1.5$ are heavily affected by the presence of outlier as are the OLS and the MDPDE with $\alpha=0.1$.
We have also performed the Wald-type tests for testing the validity of the MM model for this dataset. Interestingly, all these tests (for all $\alpha\geq 0$) reject the null hypothesis of insignificance of the MM model with a p-value $\leq 0.001$ in all cases. The outlier has no effect on the validity of MM model for this example, and hence all values of $\alpha$ lead to the same inference as the classical Wald test.
Table 8: Parameter estimates of the MM model fitted to the Drug concentration data, along with their asymptotic standard errors (in parenthesis) and 30% TAPE. [For MDPDE, $\widehat{\alpha}=0.39$ is the optimum choice obtained by IWJ method]
MDPDE
$\alpha$ 0 0.05 0.1 0.2 0.39 0.5 0.7 1
$\beta_{1}$ 89.212 88.679 88.143 86.479 80.892 80.893 80.893 80.892
(6.202) (6.174) (6.121) (5.697) (0.002) (0.006) (0.008) (0.006)
$\beta_{2}$ 0.166 0.167 0.167 0.168 0.111 0.111 0.111 0.111
(0.054) (0.054) (0.054) (0.052) (0.000) (0.000) (0.000) (0.000)
$\sigma$ 8.612 8.557 8.44 7.732 0.003 0.008 0.011 0.007
(39.646) (39.269) (38.546) (33.307) (0.000) (0.000) (0.000) (0.000)
TAPE 18.488 17.885 17.279 13.616 5.447 5.447 5.446 5.446
KPS M-est M-est
w=0.5 w=1 w=1.5 Tukey Huber
$\beta_{1}$ 132.890 132.741 131.901 82.562 82.986
(2.557) (2.967) (3.479) (1.869) (1.649)
$\beta_{2}$ 0.388 0.386 0.380 0.190 0.170
(0.031) (0.036) (0.042) (0.020) (0.017)
TAPE 1.012 0.950 0.783 1.480 4.285
Figure 5: Drug concentration data and the plot of the fitted MM models using different methods
6.3 Tyrosinase enzyme data
Both our previous examples had strong outlier effects on estimation and/or testing results. In this final example, we illustrate the performances of the proposed MDPDE based procedures as compared to its competitors for the cases where there is no outlier or only some mild outliers. We consider data from Marasovic et al. [28] which contain two sets of observations on reaction rates and concentrations of the tyrosinase ubiquitous enzyme extracted form two species of mushroom, viz Agaricus bisporus (Ab) and Pleurotus ostreatus (Po). The first set of observations on Ab contain no outlier while there is a mild outlier in the second set on Po. We fitted the MM model (28) to these two datasets using the proposed MDPDE and different competitors. Table 9 contains the parameter estimates, their standard errors and $20\%$ TAPEs; the model fits are shown in Figure 6 along with the scatter plots of the data.
The IWJ method was used for selection of optimum tuning parameter value for these datasets. For Agaricus bisporus (Ab) it took 5 iterations to converge to the optimum value $\widehat{\alpha}=0.35$, while for Pleurotus ostreatus (Po) data the method required only one iteration to select the same optimum value $\widehat{\alpha}=0.53$ for all pilot choices of $\alpha$.
Table 9: Parameter estimates of the MM model fitted to the Tyrosinase enzyme data, along with their asymptotic standard errors (in parenthesis) and $20\%$ TAPE. [For MDPDE, $\widehat{\alpha}=0.35$ and $0.53$ are the optimum choice for Ab and Po data, respectively, obtained by IWJ method]
Agaricus bisporus (Ab)
MDPDE
$\alpha$ 0 0.05 0.1 0.2 0.35 0.5 0.7 1
$\beta_{1}$ 0.377 0.377 0.376 0.376 0.378 0.378 0.378 0.376
(0.006) (0.006) (0.006) (0.006) (0.003) (0.002) (0.001) (0.003)
$\beta_{2}$ 545 545 545 545 577 588 588 578
(36.59) (36.51) (36.32) (35.53) (22.45) (9.87) (8.89) (17.38)
$\sigma$ 0.008 0.008 0.008 0.008 0.005 0.002 0.002 0.003
(0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000)
TAPE 2.88E-05 2.82E-05 2.72E-05 2.58E-05 1.26E-05 1.07E-05 1.07E-05 1.28E-05
KPS M-est M-est
w=0.5 w=1 w=1.5 Tukey Huber
$\beta_{1}$ 0.367 0.380 0.380 0.380 0.380
(0.008) (0.001) (0.001) (0.000) (0.000)
$\beta_{2}$ 519 599 599 599 598
(51.29) (3.87) (4.81) (2.14) (3.26)
TAPE 3.49E-05 8.14E-06 8.16E-06 8.03E-06 8.07E-06
Pleurotus ostreatus (Po)
MDPDE
$\alpha$ 0 0.05 0.1 0.2 0.3 0.53 0.7 1
$\beta_{1}$ 94.61 94.44 94.18 93.47 92.04 89.13 89.27 89.29
(8.683) (8.596) (8.480) (8.018) (6.642) (0.147) (0.495) (0.652)
$\beta_{2}$ 5693 5699 5699 5699 5699 5788 5809 5811
(1095) (1086) (1075) (1024) (861) (20) (67) (88)
$\sigma$ 2.277 2.250 2.210 2.059 1.672 0.034 0.111 0.135
(2.319) (2.273) (2.212) (1.976) (1.352) (0.001) (0.007) (0.011)
TAPE 1.366 1.340 1.318 1.303 1.439 0.731 0.732 0.732
KPS M-est M-est
w=0.5 w=1 w=1.5 Tukey Huber
$\beta_{1}$ 95.65 90.01 90.03 102.41 99.55
(8.015) (3.912) (4.708) (3.527) (5.892)
$\beta_{2}$ 6121 5876 5898 6938 6538
(1044) (529) (637) (464) (768)
TAPE 1.059 0.738 0.739 0.785 0.823
The scatter plot of Ab data is provided in the Figure 6(a) along with the MDPDE fits (for various $\alpha\geq 0$) and its robust competitors. The figure shows that the proposed MDPDE (for all $\alpha\geq 0$), as well as all robust competitors provide good fit. However, Figure 6(b), corresponding to the Po dataset, shows that the MDPDEs with $\alpha\geq 0.53$ and KPS with $w=1$ provide good robust fits with significantly lower TAPE, than other robust competitors.
We have also tested the validity of the MM model for these data using the Wald-type tests, which resulted a p-value of $<0.001$ for all values of $\alpha\geq 0$ (including the classical Wald test). It shows that the assumed MM model is valid and all tests give the same inference in the presence of no or mild contaminations.
(a) Agaricus bisporus (Ab)
(b) Pleurotus ostreatus (Po)
Figure 6: Tyrosinase enzyme data and the plot of the fitted MM models using different methods
7 Concluding remarks
In this paper we have proposed a robust method for parameter estimation under an NLR model based on the concept of minimum DPD estimation, introduced by Basu et al. [6]. We have established the desired asymptotic properties of our estimator and provided some mild conditions on the covariates and the regression function, under which those asymptotic properties hold. Their robustness has been verified by showing that these estimators have bounded influence functions. Under the Michaelis-Menten model in enzyme kinetics, extensive simulation study has been performed to empirically justify the robust performance of our method, which is then applied to robustly analyze several real-life datasets. Another fundamental paradigm of statistical inference, namely the hypothesis testing problem, has also been discussed in this paper. The MDPDE based Wald-type tests have been developed to test the significance of Michaelis-Menten model. The robustness of the testing procedure has also been validated through extensive simulation studies and theoretically via influence function analyses.
Our method can be treated as an optimization procedure with the minimum DPD objective function because, in most cases, minimizing the objective function is comparatively easier than solving the estimating equations. But this optimization is also a challenging problem because of the complexity of the objective function, particularly when the dimension of the regression parameter is high, where it may take a considerable amount of time. However, in our simulation experiments with the MM model, we have observed that the computation of the MDPDEs took less time than the existing robust methods. One also needs to carefully choose the initial parameter values to avoid the convergence of the optimization algorithm to a local minima. As a simple solution, one can run the algorithm for multiple initial values and finally take the solution with the lowest value of the objective function.
Finally, we would like to note that our NLR model is assumed to be homoscedastic and the errors are independently distributed. But these assumptions might not always be true in real-life situations involving NLR models. For example, some variants of the MM model (see, e.g., [33]) assume that the responses have constant coefficient of variation instead of having constant variance, implying heteroscedasticity in the assumed NLR model. So, an important future extension of our MDPDE based procedures developed in this paper would be to further extend them for such heteroscedastic NLR models as well as for the NLR models with dependent (e.g., autocorrelated) errors. We hope to pursue them in our sequel research works.
Appendix A Some background concepts
A.1 The DPD and the MDPDE
Basu et al. [6] proposed a robust method of estimation by minimizing the density power divergence (DPD) between the true data generating distribution (estimated suitably from the observed data) and the model distribution under the setup of IID data. Depending on a tuning parameter $\alpha\geq 0$, the DPD measure between two densities $f$ and $g$, defined with respect to a common dominating measure, is given by
$d_{\alpha}(g,f)=\left\{\begin{matrix}\int\left\{f^{1+\alpha}-\left(1+\frac{1}{\alpha}\right)f^{\alpha}g+\frac{1}{\alpha}g^{1+\alpha}\right\},&\mbox{if}&\alpha>0,\\ \int g\ln\left(\frac{g}{f}\right),&\mbox{if}&\alpha=0.\end{matrix}\right.$
Note that the DPD is defined as a continuous limit at $\alpha=0$, i.e., $d_{0}(g,f)=\lim_{\alpha\to 0}d_{\alpha}(g,f)$, which coincides with the Kullback-Leibler divergence. Also, the case $\alpha=1$ gives the squared $L_{2}$ distance. It has been well-established in the literature [8] that the tuning parameter $\alpha$ controls the trade-off between efficiency and robustness of the associated minimum DPD estimator (MDPDE) defined below.
Let $G$ be the (unknown) true data generating distribution with density $g$ and we want to model $g$ by the parametric family of densities =$\mathcal{F}=\left\{f_{\bm{\theta}}(\cdot):\bm{\theta}\in\bm{\Theta}\subseteq\mathbb{R}^{p}\right\}$. Now, in the minimum DPD approach, the most appropriate model density to fit $g$ well is that member of $\mathcal{F}$ for which $d_{\alpha}(g,f_{\bm{\theta}})$ is minimum over $\bm{\theta}\in\bm{\Theta}$. So, the minimum DPD functional at true distribution $G$ is defined by [6, 8]
[…]
Now, if an IID sample $x_{1},x_{2},\ldots,x_{n}$ is available from $G$, then the MDPDE $\widehat{\bm{\theta}}_{\alpha}$ of $\bm{\theta}$ can be obtained by replacing $G$ with the empirical distribution function $G_{n}$ in the above definition, leading to
$\widehat{\bm{\theta}}_{\alpha}=\bm{T}_{\alpha}(G_{n})=\arg\min_{\bm{\theta}\in\bm{\Theta}}\left\{\int f_{\bm{\theta}}^{1+\alpha}-\left(1+\frac{1}{\alpha}\right)\frac{1}{n}\sum_{i=1}^{n}f_{\bm{\theta}}^{\alpha}(x_{i})\right\}.$
Note that, unlike many other minimum divergence estimators, this DPD based method avoids the (kernel) density estimation of $g$ and the associated bandwidth selection problems even under continuous models. The consistency and asymptotic normality of the MDPDEs are proved for all $\alpha\geq 0$ under suitable conditions; see, e.g., [8], for further details and examples.
Ghosh and Basu [14] generalized the definition of the MDPDE for independent non-homogeneous (INH) observations. Suppose we have $n$ independent observations $y_{1},y_{2},\ldots,y_{n}$ with $y_{i}\sim g_{i},\mbox{ for each }i=1,2,\ldots,n$, where $g_{i}$’s are possibly different densities and we wish to model $g_{i}$ by the parametric model family $\mathcal{F}_{i}=\left\{f_{i}(\cdot;\bm{\theta}):\bm{\theta}\in\bm{\Theta}\right\}$, for each $i=1,2,\ldots,n$. Note that, in this INH setup, although the distributions may be different, all of them must depend on the same parameter $\bm{\theta}$. Ghosh and Basu [14] defined the MDPDE of $\bm{\theta}$ as the minimizer of the average DPD measure between the data points and the corresponding models, given by
$\frac{1}{n}\sum_{i=1}^{n}d_{\alpha}(\widehat{g}_{i},f_{i}(\cdot;\bm{\theta})),$
where $\widehat{g}_{i}$ is the density of the empirical distribution function of the data $y_{i}$, for each $i$. Upon simplifications, the MDPDE $\widehat{\bm{\theta}}_{\alpha}$ can equivalently be obtained by minimizing the objective function
$H_{n}(\bm{\theta})=\frac{1}{n}\sum_{i=1}^{n}V_{i}(y_{i};\bm{\theta}),$ (36)
[…]
where $\bm{u}_{i}(y;\bm{\theta})=\nabla\ln f_{i}(y;\bm{\theta})$ denotes the score function.
Note that, as $\alpha\to 0$, the MDPDE is to be obtained by minimizing $\sum_{i=1}^{n}\left[-\ln\left(f_{i}(y_{i};\bm{\theta})\right)\right]$, or equivalently maximizing $\sum_{i=1}^{n}\ln\left(f_{i}(y_{i};\bm{\theta})\right)$ with respect to $\bm{\theta}$, or by solving the estimating equation $\sum_{i=1}^{n}\bm{u}_{i}(y_{i};\bm{\theta})=0$. So, the MDPDE coincides with the MLE at $\alpha=0$.
Further, under this INH setup, the minimum DPD functional corresponding to the MDPDE $\widehat{\bm{\theta}}_{\alpha}$ at the true distributions $\bm{G}=(G_{1},G_{2},\ldots,G_{n})$, where $G_{i}$ is the distribution function associated with $g_{i}$ for each $i$, is given by
$\bm{T}_{\alpha}(\bm{G})=\arg\min_{\bm{\theta\in\bm{\Theta}}}\frac{1}{n}\sum_{i=1}^{n}d_{\alpha}(g_{i},f_{i}(\cdot;\bm{\theta})).$ (37)
Considering the contaminated density for $i$-th observation being $g_{i,\epsilon}=(1-\epsilon)g_{i}+\epsilon\wedge_{t_{i}}$, where $\wedge_{t_{i}}$ is the density of degenerate distribution at the contamination point $t_{i}$ for $i=1,2,\ldots,n$, Ghosh and Basu [14] obtained the influence function of the MDPDE functional $\bm{T}_{\alpha}$. In particular, for contamination only in the $i_{0}$-th observation, this influence function turns out to be
$IF_{i_{0}}(t_{i_{0}},\bm{T}_{\alpha},\bm{G})=\bm{\Psi}_{n}^{-1}\frac{1}{n}\left[f_{i_{0}}(t_{i_{0}};\bm{\theta}^{g})^{\alpha}\bm{u}_{i_{0}}(t_{i_{0}};\bm{\theta}^{g})-\bm{\xi}_{i_{0}}\right],$ (38)
while the same with contamination in all the observations, at contamination points in $\bm{t}=(t_{1},\ldots,t_{n})$, is given by
$IF(\bm{t},\bm{T}_{\alpha},\bm{G})=\bm{\Psi}_{n}^{-1}\frac{1}{n}\sum_{i=1}^{n}\left[f_{i}(t_{i};\bm{\theta}^{g})^{\alpha}\bm{u}_{i}(t_{i};\bm{\theta}^{g})-\bm{\xi}_{i}\right],$ (39)
[…]
A.2 Existing robust estimators under the NLR models
M-estimators:
For the NLR model given in (3), an M-estimator $\widehat{\bm{\beta}}_{M}$ of $\bm{\beta}$ is defined as
$\widehat{\bm{\beta}}_{M}=\arg\min_{\bm{\Theta_{\beta}}}\sum_{i=1}^{n}\rho\left(\frac{r_{i}}{\sigma}\right),$ (40)
where $r_{i}=y_{i}-\mu(\bm{x_{i}},\bm{\beta}),\mbox{ for }i=1,\ldots,n$, $\bm{\Theta_{\beta}}$ is the parameter space of the regression parameter $\bm{\beta}$ and, if $\sigma$ is unknown, it needs to be estimated by any usual robust scale estimate, e.g., $\widehat{\sigma}=1.4826\mbox{Med}_{i}(|r_{i}-\mbox{Med}_{j}r_{j}|)$, and $\rho:\mathbb{R}\to\mathbb{R}^{+}$ is an appropriate even function which is non-decreasing in $\mathbb{R}^{+}$. When $\rho(\cdot)$ is differentiable and $\psi=\rho^{{}^{\prime}}$ then the associated estimating equation is given by
$\sum_{i=1}^{n}\psi\left(\frac{y_{i}-\mu(\bm{x_{i}},\bm{\beta})}{\sigma}\right)\nabla_{\bm{\beta}}\mu(\bm{x_{i}},\bm{\beta})=0.$ (41)
Following the asymptotic normality results of regression M estimators from Maronna et al. [29], it can be shown that under certain conditions the M-estimator $\bm{\widehat{\beta}}_{M}$, obtained by solving (41), satisfies
[…]
where $v_{M}=\sigma^{2}\frac{E\psi^{2}(\frac{u}{\sigma})}{\left[E\psi^{{}^{\prime}}(\frac{u}{\sigma})\right]^{2}},\mbox{ with }u\sim\mathcal{N}(0,\sigma^{2})$. This asymptotic variance formula has been used to compute the ARE for the M-estimator of $\bm{\beta}$, presented in Section 2.2.
Different forms of the function $\rho(\cdot)$ lead to different M-estimators. Liu et al. [27] considered the Huber’s $\rho$ function (provided by Huber [19]) given by
$\rho_{c}(x)=\left\{\begin{matrix}\frac{x^{2}}{2},&\mbox{if}&|x|\leq c,\\ cx-\frac{c^{2}}{2},&\mbox{if}&|x|>c,\end{matrix}\right.$
with usual choice $c=1.345$ for analysing different NLR models. Later, Marasovic et al. [28] studied M-estimators only for the MM model (and not for general NLR models) with the Tukey’s biweight $\rho$ function, given by
$\rho_{c}(x)=\left\{\begin{matrix}\frac{c^{2}}{6}(1-[1-(\frac{x}{c})^{2}]^{3}),&\mbox{if}&|x|\leq c,\\ \frac{c^{2}}{6},&\mbox{if}&|x|>c.\end{matrix}\right.$
In this case, the tuning parameter $c=4.685$ yields $95\%$ efficiency under normal error assumption and can successfully handle contamination up to $10\%$ outliers (as illustrated in Section 2.1 of [28]).
The KPS estimator:
Tabatabai et al. [36] proposed a robust estimator of $\bm{\beta}$ under the NLR model in (3), by suitably modifying the M-estimation approach, as given by
$\bm{\widehat{\beta}}=\arg\min_{\bm{\beta}\in\mathbb{R}^{p}}\sum_{i=1}^{n}\frac{\rho_{\omega}(t_{i})}{L_{i}},$
where $\rho_{\omega}(x)=1-Sech(\omega x)$ with $Sech(x)=\frac{2}{e^{x}+e^{-x}}$, $t_{i}=\frac{1}{\sigma}(1-h_{ii})(y_{i}-\mu(\bm{x_{i}},\bm{\beta}))$, with $h_{ii}$ being the $i$-th diagonal element of the matrix $\bm{H}=\bm{\dot{\mu}(\dot{\mu}^{T}\dot{\mu})^{-1}\dot{\mu}^{T}}$ and $L_{i}=\sum_{j=1}^{k}\max\{M_{j},|x_{ij}|\},\ i=1,\ldots,n$, where $M_{j}=\mbox{Median}\{|x_{1j}|,\ldots,|x_{nj}|\},\ j=1,\ldots,k$, with $x_{ij}$ be the $j$-th element of the $k$-dimensional covariate vector $\bm{x_{i}}$ of the $i$-th observation.
If $\sigma$ is unknown, they suggested to use either of the two robust estimates defined as $\widehat{\sigma}=1.1926\mbox{ Med}_{1\leq i\leq n}\left(\mbox{Med}_{1\leq j\leq n}|r_{i}-r_{j}|\right)$ or $\widehat{\sigma}=2.2219\ \left\{|r_{i}-r_{j}|:1\leq i<j\leq n\right\}_{(l)}$, with $r_{i}=y_{i}-\mu\left(\bm{x_{i}},\bm{\widehat{\beta}}\right)$, $l=\dbinom{\left[\frac{n}{2}\right]+1}{2}$, the binomial coefficient, and $S_{(l)}$ denotes the $l$-th order statistic in the set $S$.
The MOM-estimator:
Liu et al. [26] proposed a very simple robust estimation method for the NLR model (3), called the median of means (MOM) method. This method consists of the following steps:
  • Dataset of size $n$ is divided randomly into $g$ groups of equal size, assuming $n$ is divisible by $g$. Some suitable suggestion for the choice of g is mentioned in [26].
  • Obtain the OLS estimate of $\bm{\beta}$ from every group of observations separately, which we denote as $\widehat{\bm{\beta}}^{(j)}=\left(\widehat{\beta}_{1}^{(j)},\ldots,\widehat{\beta}_{p}^{(j)}\right)^{T}$ for group $j,\ j=1,2,\ldots,g$.
  • The MOM estimator is then given by $\widehat{\bm{\beta}}^{MOM}=\left(\widehat{\beta}_{1}^{MOM},\ldots,\widehat{\beta}_{p}^{MOM}\right)^{T}$, where $\widehat{\beta}_{k}^{MOM}=\mbox{Median}\left\{\widehat{\beta}_{k}^{(1)},\ldots,\widehat{\beta}_{k}^{(g)}\right\},\mbox{ for all }k=1,2,\ldots,p$.
Liu et al. [26] have also proved asymptotic properties (consistency and asymptotic normality) of the estimator under certain assumptions. However, from the definition of the method, it is intuitively clear that this method gives good, stable performance only when the sample size is moderately large and the proportion of outlier is small, which is also justified through our simulation studies, presented in the Section 4.3.
Appendix B Proof of theorems
B.1 Proof of Theorem 2.1
First, we will state and prove a lemma and then use Theorem 3.1 of Ghosh and Basu [14] to complete the proof of our theorem.
Lemma B.1.
Consider the setup of the NLR model (3) and suppose that the true data generating distributions belong to the assumed parametric model families. Then Conditions (R1) and (R2) imply Assumptions (A1)-(A7) of Ghosh and Basu [14].
Proof.
As the true distributions belong to the corresponding parametric model families, under which the distribution of $y_{i}$ is $\mathcal{N}\left(\mu(\bm{x_{i}},\bm{\beta}),\sigma^{2}\right)$ with $\mu(\bm{x_{i}},\bm{\beta})$ being thrice continuously differentiable, Assumptions (A1)-(A3) of Ghosh and Basu [14] directly follow from the property of the normal density in view of (R1). Assumption (A4) follows from Equation (14) of Condition (R2). The form of $V_{i}(y_{i};\bm{\theta})$ in (4) ensures that Assumption (A5) follows from Equation (13) of (R1). Now, to prove Equation (3.8) of Assumption (A6) of Ghosh and Basu [14], we note that
$\nabla_{j}V_{i}(y_{i};\bm{\theta})=\kappa e^{-\frac{\alpha}{2\sigma^{2}}(y_{i}-\mu(\bm{x_{i}},\bm{\beta}))^{2}}(y_{i}-\mu(\bm{x_{i}},\bm{\beta}))(\bm{\dot{\mu}})_{ij},~{}~{}~{}~{}j=1,\ldots,p,$
where $\nabla_{j}$ is the first order partial derivative with respect to the $j$-th component of $\bm{\beta}$ and $\kappa=-\frac{1+\alpha}{(2\pi)^{\alpha/2}\sigma^{\alpha+2}}$. Then,
$\frac{1}{n}\sum_{i=1}^{n}E_{i}[|\nabla_{j}V_{i}(y_{i};\bm{\theta})|I(|\nabla_{j}V_{i}(y_{i};\bm{\theta})|>N)]$
$=|\kappa|\frac{1}{n}\sum_{i=1}^{n}E_{i}\left[e^{-\frac{\alpha z_{i}^{2}}{2\sigma^{2}}}|z_{i}||(\bm{\dot{\mu}})_{ij}|\times I\left(e^{-\frac{\alpha z_{i}^{2}}{2\sigma^{2}}}|z_{i}||(\bm{\dot{\mu}})_{ij}|>\frac{N}{|\kappa|}\right)\right],\mbox{ letting }z_{i}=y_{i}-\mu(\bm{x_{i}},\bm{\beta})$
$\leq|\kappa|\frac{1}{n}\sum_{i=1}^{n}|(\bm{\dot{\mu}})_{ij}|E_{i}\left[e^{-\frac{\alpha z_{i}^{2}}{2\sigma^{2}}}|z_{i}|\times I\left(e^{-\frac{\alpha z_{i}^{2}}{2\sigma^{2}}}|z_{i}|>\frac{N}{|\kappa|(\sup_{n>1}\max_{1\leq i\leq n}|(\bm{\dot{\mu}})_{ij}|)}\right)\right]$
$=|\kappa|E_{1}\left[e^{-\frac{\alpha z_{1}^{2}}{2\sigma^{2}}}|z_{1}|\times I\left(e^{-\frac{\alpha z_{1}^{2}}{2\sigma^{2}}}|z_{1}|>\frac{N}{|\kappa|(\sup_{n>1}\max_{1\leq i\leq n}|(\bm{\dot{\mu}})_{ij}|)}\right)\right]\left(\frac{1}{n}\sum_{i=1}^{n}|(\bm{\dot{\mu}})_{ij}|\right),$
as $z_{i}$’s are IID. Since we have $\sup_{n>1}\max_{1\leq i\leq n}|(\bm{\dot{\mu}})_{ij}|=O(1)$ by (12) of (R1), so, by dominated convergence theorem (DCT)
$\lim_{N\to\infty}E_{1}\left[e^{-\frac{\alpha z_{1}^{2}}{2\sigma^{2}}}|z_{1}|\times I\left(e^{-\frac{\alpha z_{1}^{2}}{2\sigma^{2}}}|z_{1}|>\frac{N}{|\kappa|(\sup_{n>1}\max_{1\leq i\leq n}|(\bm{\dot{\mu}})_{ij}|)}\right)\right]=0.$
We also have
$\sup_{n>1}\left(\frac{1}{n}\sum_{i=1}^{n}|(\bm{\dot{\mu}})_{ij}|\right)\leq\sup_{n>1}\max_{1\leq i\leq n}|(\bm{\dot{\mu}})_{ij}|=O(1).$
Hence, Equation (3.8) of (A6) follows for all $j=1,2,\ldots,p$. Similarly one can show that it follows also for $j=p+1$, i.e., for $\nabla_{\sigma^{2}}$, the gradient with respect to $\sigma^{2}$. Similarly, Equation (3.9) of Assumption (A6) and Assumption (A7) of Ghosh and Basu [14] can also be shown to hold using second part of (12) of (R1) and (15) of (R2), respectively. ∎
Proof of the theorem:
The consistency part follows from Lemma B.1 and Theorem 3.1 of Ghosh and Basu [14]. Also, using Lemma (B.1), it follows from Theorem 3.1 of Ghosh and Basu [14] that the asymptotic distribution of $\bm{\Omega}_{n}^{-\frac{1}{2}}\bm{\Psi}_{n}\left[\sqrt{n}(\bm{\widehat{\theta}_{\alpha}}-\bm{\theta}_{0})\right]$ is $\mathcal{N}_{p+1}(\bm{0},\mathbb{I}_{p+1})$, where $\bm{\Omega}_{n}$ and $\bm{\Psi}_{n}$ are as defined in Equations (3.3) and (3.4) in Ghosh and Basu [14], and are to be computed at the model density with $\bm{\theta}=\bm{\theta_{0}}$. For the present case of NLR model (3), we can see that
$\bm{u}_{i}(y;\bm{\theta})=\nabla_{\bm{\theta}}\ln f_{i}(y;\bm{\theta})=\begin{pmatrix}\frac{(y-\mu(\bm{x_{i}},\bm{\beta}))}{\sigma^{2}}\nabla_{\bm{\beta}}\mu(\bm{x_{i}},\bm{\beta})\\ \\ \frac{(y-\mu(\bm{x_{i}},\bm{\beta}))^{2}}{2\sigma^{4}}-\frac{1}{2\sigma^{2}}\end{pmatrix}_{(p+1)\times 1}.$ (42)
After some simple calculations we obtain the $(p+1)\times(p+1)$ matrix $\bm{J}_{\alpha}^{(i)}$ as
[…]
So, a simplified form of the matrix $\bm{\Psi}_{n}$ is given by
$\bm{\Psi}_{n}=\frac{1}{n}\sum_{i=1}^{n}\bm{J}_{\alpha}^{(i)}=\begin{bmatrix}\frac{\zeta_{\alpha}}{n}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)&\bm{0}_{p}\\ \\ \bm{0}_{p}^{T}&\varsigma_{\alpha}\end{bmatrix}.$
We also have
[…]
and using it we obtain a simplified form of the matrix $\bm{\Omega}_{n}$ as
$\bm{\Omega}_{n}=\frac{1}{n}\sum_{i=1}^{n}(\bm{J}^{(i)}_{2\alpha}-\bm{\xi}_{i}\bm{\xi}_{i}^{T})=\begin{bmatrix}\frac{\zeta_{2\alpha}}{n}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)&\bm{0}_{p}\\ \\ \bm{0}_{p}^{T}&\varsigma_{2\alpha}-\frac{\alpha^{2}}{4}\zeta_{\alpha}^{2}\end{bmatrix}.$
Now using these simplified form of $\bm{\Psi}_{n}$ and $\bm{\Omega}_{n}$, at $\bm{\theta}=\bm{\theta}_{0}$, we have
$\bm{\Omega}_{n}^{-\frac{1}{2}}\bm{\Psi}_{n}\left[\sqrt{n}\left(\bm{\widehat{\theta}_{\alpha}}-\bm{\theta_{0}}\right)\right]=\begin{pmatrix}\frac{\zeta_{\alpha}}{\sqrt{\zeta_{2\alpha}}}\left(\bm{\dot{\mu}(\beta_{0})}^{T}\bm{\dot{\mu}(\beta_{0})}\right)^{1/2}\left(\bm{\widehat{\beta}_{\alpha}}-\bm{\beta_{0}}\right)\\ \\ \frac{\varsigma_{\alpha}}{\sqrt{\varsigma_{2\alpha}-\frac{\alpha^{2}}{4}\zeta_{\alpha}^{2}}}\sqrt{n}\left(\widehat{\sigma}_{\alpha}^{2}-\sigma_{0}^{2}\right)\end{pmatrix},$
which asymptotically follows $\mathcal{N}_{p+1}(\bm{0},\mathbb{I}_{p+1})$, where $\zeta_{p}$’s and $\varsigma_{p}$’s are to be computed at $\sigma=\sigma_{0}$. Therefore, asymptotically, $\frac{\zeta_{\alpha}}{\sqrt{\zeta_{2\alpha}}}\left(\bm{\dot{\mu}(\beta_{0})}^{T}\bm{\dot{\mu}(\beta_{0})}\right)^{1/2}\left(\bm{\widehat{\beta}_{\alpha}}-\bm{\beta_{0}}\right)$ and
$\frac{\varsigma_{\alpha}}{\sqrt{\varsigma_{2\alpha}-\frac{\alpha^{2}}{4}\zeta_{\alpha}^{2}}}\sqrt{n}\left(\widehat{\sigma}^{2}-\sigma_{0}^{2}\right)$ are independently distributed with their respective asymptotic distributions being $\mathcal{N}_{p}(\bm{0,\mathbb{I}_{p}})$ and $\mathcal{N}(0,1)$. This completes the proof.
B.2 Proof of Proposition 1
For the case of the MM model in (28), $\bm{\dot{\mu}(\beta)}$ (given in (29)) is of order $n\times 2$ and
$\sup_{n>1}\max_{1\leq i\leq n}|(\bm{\dot{\mu}})_{i1}|=\sup_{n>1}\max_{1\leq i\leq n}\frac{x_{i}}{\beta_{2}+x_{i}},$ (45)
$\sup_{n>1}\max_{1\leq i\leq n}|(\bm{\dot{\mu}})_{i2}|=|\beta_{1}|\sup_{n>1}\max_{1\leq i\leq n}\frac{x_{i}}{(\beta_{2}+x_{i})^{2}},$ (46)
$\sup_{n>1}\max_{1\leq i\leq n}|(\bm{\dot{\mu}})_{i1}(\bm{\dot{\mu}})_{i2}|=|\beta_{1}|\sup_{n>1}\max_{1\leq i\leq n}\frac{x_{i}^{2}}{(\beta_{2}+x_{i})^{3}}.$ (47)
We have $x_{i}>0$ for all $i$ and $\beta_{2}\geq 0$. So, $\frac{x_{i}}{\beta_{2}+x_{i}}$ is bounded between 00 and $1$. So, the quantity in Equation (45) is $O(1)$.
Now consider the functions $u_{1}(x)=\frac{x}{(\beta_{2}+x)^{2}}$ and $u_{2}(x)=\frac{x^{2}}{(\beta_{2}+x)^{3}}$, both defined for $x>0$. So, both $u_{1}$ and $u_{2}$ are positive real valued functions. And, by simple calculations, we get
$\frac{d}{dx}u_{1}(x)=\frac{\beta_{2}-x}{(\beta_{2}+x)^{3}}=\left\{\begin{matrix}>0,&\mbox{if}&x<\beta_{2},\\ =0,&\mbox{if}&x=\beta_{2},\\ <0,&\mbox{if}&x>\beta_{2},\end{matrix}\right.$
and
$\frac{d}{dx}u_{2}(x)=\frac{x(2\beta_{2}-x)}{(\beta_{2}+x)^{4}}=\left\{\begin{matrix}>0,&\mbox{if}&x<2\beta_{2},\\ =0,&\mbox{if}&x=2\beta_{2},\\ <0,&\mbox{if}&x>2\beta_{2}.\end{matrix}\right.$
So, $u_{1}(x)$ attains its unique maxima at $x=\beta_{2}$, and hence it is bounded in $(0,u_{1}(\beta_{2})]$. This shows that the quantity in Equation (46) is $O(1)$. Similarly, $u_{2}(x)$ attains its unique maxima at $x=2\beta_{2}$, and hence it is bounded in $(0,u_{2}(2\beta_{2})]$, which indicates that the quantity in Equation (47) is $O(1)$. Therefore, (R1) is satisfied.
B.3 Proof of Proposition 2
Let us define $A_{n}=\frac{1}{n}\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}$. The characteristic equation of $A_{n}$ is
$\lambda^{2}-tr(A_{n})\lambda+det(A_{n})=0.$
So, the eigenvalues of $A_{n}$ are given by
$\lambda=\frac{tr(A_{n})\pm\sqrt{tr(A_{n})^{2}-4det(A_{n})}}{2}.$
Here $\lambda$’s are real as $A_{n}$ is real symmetric matrix. Now, in order to have the minimum eigenvalue of $A_{n}$ strictly positive, we must have
$tr(A_{n})-\sqrt{tr(A_{n})^{2}-4det(A_{n})}>0$
$\Leftrightarrow det(A_{n})>0$
$\Leftrightarrow\sum_{i=1}^{n}a_{i}^{2}\ \ \sum_{i=1}^{n}b_{i}^{2}>\left(\sum_{i=1}^{n}a_{i}b_{i}\right)^{2},$
with $a_{i}=\frac{x_{i}}{\beta_{2}+x_{i}},b_{i}=\frac{\beta_{1}x_{i}}{(\beta_{2}+x_{i})^{2}}$, which is true by Cauchy-Schwartz inequality unless the equality occurs. The equality occurs only when $a_{i}\propto b_{i}\ \mbox{ for all }i$, i.e., when $x_{i}=$ constant, for all $i$ or $\beta_{1}=0$.
Therefore, in order to have the minimum eigenvalue of $A_{n}$ strictly positive and satisfy (14), it is necessary that the values $x_{i}$’s are not all same, for $n\geq 2$, and $\beta_{1}\neq 0$.
Appendix C Additional numerical results
C.1 Influence functions
[…]
(a) IF of $T_{\alpha}^{\beta_{1}}$ with $i_{0}=10$
(b) IF of $T_{\alpha}^{\beta_{2}}$ with $i_{0}=10$
(c) IF of $T_{\alpha}^{\beta_{1}}$ with $i_{0}=40$
(d) IF of $T_{\alpha}^{\beta_{2}}$ with $i_{0}=40$
Figure 7: IFs of the MDPDEs of the parameters $\beta_{1}$ and $\beta_{2}$ in the MM model, for contamination only in $i_{0}$-th observation under Setup 1
(a) IF of $T_{\alpha}^{\beta_{1}}$ with $i_{0}=15$
(b) IF of $T_{\alpha}^{\beta_{2}}$ with $i_{0}=15$
(c) IF of $T_{\alpha}^{\beta_{1}}$ with $i_{0}=45$
(d) IF of $T_{\alpha}^{\beta_{2}}$ with $i_{0}=45$
Figure 8: IFs of the MDPDEs of the parameters $\beta_{1}$ and $\beta_{2}$ in the MM model, for contamination only in $i_{0}$-th observation under Setup 2
C.2 Additional tables from simulation results
Table 10 presents the average prediction errors for the fits obtained by the MDPDE and its robust competitors under the MM model (28), based on 1000 replications for contaminated data with contamination proportions $e_{c}=20\%$ and $30\%$; associated EBias and EMSE of the parameter estimates are given in Tables 11 and 12, respectively.
Table 10: Average prediction errors obtained by different estimators and different amounts of contamination (the minimum values obtained for each parameter are highlighted in bold font)
Setup Setup 1 Setup 2
Outlier direction response response + covariate response response + covariate
Outlier Prop 20% 30% 20% 30% 20% 30% 20% 30%
OLS 15.46 32.75 8.214 11.59 79.51 170.43 100.69 172.67
KPS(w=0.5) 1.094 13.02 0.976 2.009 3.802 3.997 3.810 3.901
KPS(w=1) 1.041 2.745 1.080 2.258 3.827 3.800 3.829 3.809
KPS(w=1.5) 1.208 1.642 1.010 1.214 3.889 3.838 3.891 3.843
Huber 1.221 2.408 2.784 9.485 5.008 11.15 14.18 127.95
Tukey 0.947 0.941 0.947 1.829 5.930 5.548 14.81 124.12
MDPDE($\alpha$=0.05) 12.93 30.06 9.498 13.37 68.41 158.58 88.67 165.00
MDPDE($\alpha$=0.1) 10.37 27.32 8.156 11.74 54.83 162.34 80.34 181.05
MDPDE($\alpha$=0.2) 1.500 21.77 3.735 9.261 3.830 19.97 5.486 96.12
MDPDE($\alpha$=0.3) 0.953 4.785 0.950 5.183 7.194 4.564 3.800 8.863
MDPDE($\alpha$=0.5) 0.956 0.950 0.954 0.946 3.811 3.794 3.816 3.807
MDPDE($\alpha$=0.7) 0.962 0.951 0.961 0.949 3.829 3.808 3.834 3.820
MDPDE($\alpha$=1) 0.972 0.956 0.971 0.955 3.856 3.826 3.862 3.838
Table 11: Empirical absolute bias and empirical MSE obtained by different estimators in simulations with Setup 1 and different amounts of contamination (the minimum values obtained for each parameter are highlighted in bold font)
Outlier Direction Contamination only in response Contamination in both response and covariate
Outlier proportion 10% 40% 10% 40%
$\beta_{1}$ $\beta_{2}$ $\sigma$ $\beta_{1}$ $\beta_{2}$ $\sigma$ $\beta_{1}$ $\beta_{2}$ $\sigma$ $\beta_{1}$ $\beta_{2}$ $\sigma$
Empirical Bias
OLS 4.9733 2.6438 7.2248 2.7391 56.40 260.74 42.21 163.89
MOM 4.1274 3.9194 124.75 7.3124 3.1807 221.13 5.1049 356.55
KPS(w=0.5) 0.0132 0.0094 0.0221 0.1983 0.0521 0.1183 1.1789 4.7107
KPS(w=1) 0.0234 0.1011 0.0115 0.0457 0.0287 0.1117 0.0314 0.0955
KPS(w=1.5) 0.0388 0.1583 0.0247 0.1038 0.0405 0.1605 0.0335 0.1268
M-Est (Huber) 0.5502 0.0696 1.3424 0.3195 3.1837 8.8381 5.7900 9.0000
M-Est (Tukey) 0.0192 0.0855 0.0248 0.0710 0.0255 0.1010 0.5843 1.0422
MDPDE($\alpha$=0.05) 2.9360 0.5796 5.0024 5.5897 0.9690 7.4614 29.49 123.06 2.9775 32.75 116.42 4.2146
MDPDE($\alpha$=0.1) 0.0941 0.0249 0.2047 0.5167 0.9748 1.6376 1.6620 7.5578 0.1757 8.1187 31.41 0.9716
MDPDE($\alpha$=0.2) 0.0123 0.0701 0.0049 0.0103 0.0496 0.0337 0.0225 0.0977 0.0071 0.0765 0.3353 0.0347
MDPDE($\alpha$=0.3) 0.0168 0.0894 0.0171 0.0125 0.0746 0.0505 0.0222 0.1036 0.0186 0.0247 0.1083 0.0529
MDPDE($\alpha$=0.5) 0.0228 0.1145 0.0475 0.0198 0.1028 0.1097 0.0254 0.1201 0.0485 0.0261 0.1201 0.1116
MDPDE($\alpha$=0.7) 0.0272 0.1313 0.0774 0.0243 0.1214 0.1704 0.0290 0.1348 0.0783 0.0285 0.1315 0.1721
MDPDE($\alpha$=1) 0.0335 0.1537 0.1156 0.0291 0.1411 0.2499 0.0342 0.1539 0.1162 0.0318 0.1460 0.2515
Empirical MSE
OLS 36.43 99.85 83.95 192.01 7559 175837 1866 29046
MOM 49.79 612.19 6.8E+06 1737.08 12.67 136261 55.56 5.7E+06
KPS(w=0.5) 0.0712 0.5051 0.1184 0.8523 0.0700 0.4974 50.62 1037
KPS(w=1) 0.0757 0.5376 0.0864 0.6142 0.0740 0.5260 0.0822 0.5970
KPS(w=1.5) 0.0941 0.7241 0.0971 0.7193 0.0943 0.7115 0.0935 0.6786
M-Est (Huber) 0.3977 0.5981 2.4094 2.2897 10.40 79.48 34.58 81.00
M-Est (Tukey) 0.0579 0.4184 0.0699 0.5068 0.0569 0.4100 3.5212 8.6970
MDPDE($\alpha$=0.05) 18.62 96.30 32.87 40.39 23.60 60.56 1150 20448 11.73 1175 15229 19.55
MDPDE($\alpha$=0.1) 0.3157 0.6118 1.0584 26.06 266.19 13.62 58.44 1237 0.5567 296.85 4527 4.2317
MDPDE($\alpha$=0.2) 0.0612 0.4511 0.0148 0.0818 0.5559 0.0578 0.0582 0.4177 0.0151 2.5361 54.08 0.0307
MDPDE($\alpha$=0.3) 0.0623 0.4568 0.0158 0.0743 0.5525 0.0221 0.0601 0.4336 0.0160 0.0707 0.5117 0.0226
MDPDE($\alpha$=0.5) 0.0666 0.4946 0.0209 0.0772 0.5721 0.0355 0.0648 0.4791 0.0210 0.0748 0.5493 0.0362
MDPDE($\alpha$=0.7) 0.0716 0.5481 0.0280 0.0808 0.6101 0.0570 0.0701 0.5346 0.0282 0.0792 0.5952 0.0578
MDPDE($\alpha$=1) 0.0793 0.6320 0.0402 0.0863 0.6760 0.0965 0.0784 0.6238 0.0404 0.0848 0.6623 0.0976
Table 12: Empirical absolute bias and empirical MSE obtained by different estimators in simulations with Setup 2 and different amounts of contamination (the minimum values obtained for each parameter are highlighted in bold font)
Outlier Direction Contamination only in response Contamination in both response and covariate
Outlier proportion 10% 40% 10% 40%
$\beta_{1}$ $\beta_{2}$ $\sigma$ $\beta_{1}$ $\beta_{2}$ $\sigma$ $\beta_{1}$ $\beta_{2}$ $\sigma$ $\beta_{1}$ $\beta_{2}$ $\sigma$
Empirical Bias
OLS 10.54 0.2015 15.72 0.24382 28.55 7.6917 41.55 11.09
MOM 8.3100 0.4567 15.99 0.7324 28.70 6.6655 47.51 11.22
KPS(w=0.5) 0.0666 0.0289 0.2759 0.1323 0.0822 0.0411 0.1605 0.1090
KPS(w=1) 0.0003 0.0016 0.0134 0.0109 0.0180 0.0094 0.0575 0.0309
KPS(w=1.5) 0.0186 0.0042 0.0198 0.0014 0.0050 0.0011 0.0125 0.0117
M-Est (Huber) 0.8973 0.0919 2.2299 0.1811 4.8576 1.0918 27.05 5.8772
M-Est (Tukey) 0.0685 0.0031 0.2401 0.0128 3.5690 0.9716 26.05 5.7811
MDPDE($\alpha$=0.05) 0.5644 0.5019 5.3485 0.3564 1.2477 12.89 7.3678 1.3647 7.1999 16.44 3.5317 14.66
MDPDE($\alpha$=0.1) 0.0341 0.0288 0.0947 0.2141 0.0747 0.7250 0.0764 0.0528 0.1385 1.7809 0.2961 1.4378
MDPDE($\alpha$=0.2) 0.0119 0.0059 0.0054 0.0125 0.0108 0.0601 0.0422 0.0199 0.0225 0.0797 0.0419 0.0971
MDPDE($\alpha$=0.3) 0.0033 0.0022 0.0277 0.0051 0.0036 0.1037 0.0251 0.0120 0.0368 0.0419 0.0244 0.1232
MDPDE($\alpha$=0.5) 0.0014 0.0003 0.0840 0.0142 0.0000 0.2200 0.0131 0.0067 0.0873 0.0166 0.0132 0.2280
MDPDE($\alpha$=0.7) 0.0039 0.0007 0.1416 0.0191 0.0019 0.3398 0.0076 0.0043 0.1418 0.0077 0.0094 0.3419
MDPDE($\alpha$=1) 0.0071 0.0018 0.2163 0.0219 0.0029 0.4972 0.0029 0.0025 0.2136 0.0000 0.0063 0.4942
Empirical MSE
OLS 127.86 1.2657 268.46 1.3927 1026.58 104.58 1993.45 181.41
MOM 154.28 3.5655 383.09 6.8593 1324.69 119.51 2784.57 227.17
KPS(w=0.5) 0.4331 0.0375 0.8790 0.1089 0.4197 0.0382 0.5939 0.0709
KPS(w=1) 0.4821 0.0392 0.5435 0.0437 0.4766 0.0391 0.5443 0.0469
KPS(w=1.5) 0.6142 0.0497 0.6296 0.0495 0.6076 0.0495 0.6297 0.0518
M-Est (Huber) 1.7325 0.1060 12.83 0.4652 91.42 6.7782 912.82 45.14
M-Est (Tukey) 2.9843 0.0343 9.4442 0.2652 104.27 7.8639 892.73 44.69
MDPDE($\alpha$=0.05) 63.26 11.48 102.59 151.63 8.7679 365.28 294.77 49.97 116.81 918.56 83.04 370.88
MDPDE($\alpha$=0.1) 0.4434 0.0463 1.0514 5.3745 0.2249 12.67 0.7096 0.0522 0.8664 65.63 4.1487 22.13
MDPDE($\alpha$=0.2) 0.4017 0.0345 0.0593 0.4890 0.0428 0.0828 0.4047 0.0357 0.0641 0.5229 0.0522 0.1032
MDPDE($\alpha$=0.3) 0.4112 0.0350 0.0638 0.4857 0.0409 0.0909 0.4146 0.0358 0.0663 0.5135 0.0470 0.0996
MDPDE($\alpha$=0.5) 0.4451 0.0381 0.0828 0.5136 0.0432 0.1449 0.4473 0.0385 0.0839 0.5264 0.0460 0.1489
MDPDE($\alpha$=0.7) 0.4821 0.0415 0.1088 0.5412 0.0455 0.2289 0.4848 0.0419 0.1090 0.5536 0.0481 0.2300
MDPDE($\alpha$=1) 0.5374 0.0463 0.1521 0.5806 0.0489 0.3834 0.5420 0.0470 0.1517 0.5934 0.0512 0.3794